### v1

# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

Abstract
The first production release of the CUDACPP plugin for the Madgraph5_aMC@NLO generator, which speeds up matrix element (ME) calculations for leading-order (LO) QCD processes using a data parallel approach on vector CPUs and GPUs, was delivered in October 2024. This has been described in previous publications by the team behind that effort. In this paper, I describe my work on some additional developments providing further optimizations of CUDACPP for GPUs, which I consider ready for inclusion in a new release of the software. The new approach mainly consists in splitting the calculation of the ME, which has been so far performed using a single large GPU kernel, into several smaller kernels. I also take this opportunity to describe more in detail some features of the CUDACPP software that are relevant to these new developments and that have not yet been documented.
journal: Eur. Phys. J. Cjournal: Eur. Phys. J. C
1 Introduction
The MadGraph5_aMC@NLO bib:mg5 ; bib:mg5amc physics event generator (in the following, MG5aMC) is an essential component of the data processing and analysis workflows of many high-energy physics (HEP) experiments, notably those at CERN’s Large Hadron Collider (LHC). Using Monte Carlo (MC) techniques, MG5aMC allows the calculation of cross sections and the generation of events to provide theoretical predictions against which experimental measurements can be compared. The MG5aMC software has been developed over more than two decades and does not yet fully exploit the potential of modern computing architectures. As is the case for other generators, this is a source of concern because event generation in the LHC experiments has a large computational cost, which is predicted to further increase during the High-Luminosity LHC (HL-LHC) programme bib:csbs .
In this context, a recent progress has been the port and optimization of the MG5aMC software for data parallel processing on graphical processing units (GPUs) and on CPUs with SIMD (Single Instruction Multiple Data) vector instructions. After a five-year development programme, which was initially facilitated by the activities of the HEP Software Foundation (HSF) event generator working group bib:csbs , and which regularly reported its progress through conference proceedings bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 , this project delivered its first production release of the MG5aMC code with GPU and SIMD support for QCD leading-order (LO) processes in October 2024. This has been reported at the CHEP2024 conference bib:chep2024 and in a more recent journal submission bib:zw2025 . MG5aMC is a code-generating framework, largely written in Python, which allows users to generate code, by default using Fortran, for any physics process of their choice. The computational bottleneck of the MG5aMC workflow is the calculation of the matrix element (ME) for each phase space point from the momenta of the initial and final state particles in that event. The main outcome of this project is a new code-generating plugin, which was named “CUDACPP” because the generated physics code uses C++ instead of Fortran to execute the calculation of MEs on vector CPUs, and also includes CUDA extensions to run the code on NVidia GPUs (as well as HIP extensions for AMD GPUs). The new code speeds up the ME computation by performing it for many events at the same time using a data parallel approach bib:hsf2020 ; bib:lhcc2020 . To achieve this, work was also needed in the process-agnostic core MG5aMC event processing algorithm for LO processes, MadEvent, which is written in Fortran and with which the process-specific C++ code generated by CUDACPP is interfaced and linked: the main change is the move from a sequential single-event API to a parallel multi-event API, and the refactoring of some functions to make them stateless and re-entrant. This effort has been the result of a collaboration involving many existing and new contributors of the MG5aMC project, as detailed in the references cited above bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 .
In this article, I describe a few new enhancements of the CUDACPP plugin, specifically targeting more efficient processing on GPUs. These are developments for which the initial brainstorming with some of my colleagues from the development team, notably Olivier Mattelaer who prototyped with me the first concrete changes to the code, took place during the CSCS GPU Hackathon that we attended in Lugano in September 2022 bib:hack2022 . I then invested further effort in this area in February and especially in October 2024 bib:ks2024 after the first CUDACPP release, but had to pause this work again due to other constraints. In September 2025, I was finally able to resume this activity, and I believe that some of these developments are now ready to be considered for their inclusion in a new production release, which is why I found it useful to document them in some detail. The main idea behind all of these developments can be referred to as “kernel splitting”. In short, this consists in replacing the large monolithic kernel currently used by CUDACPP for the calculation of MEs from particle momenta, named sigmaKin, by many smaller kernels, executed either in parallel or sequentially, to achieve a more efficient and scalable use of the GPU. The aim is not only that of possibly achieving a higher peak throughput (in terms of MEs computed per second) with a large GPU grid, i.e. with a large number of events processed in parallel on the GPU during one offloading cycle, but also that of achieving a faster increase of throughput already with smaller GPU grids. An important practical difference between the traditional MG5aMC workflow on CPUs and that using GPUs, in fact, is that the former was designed to handle iterations involving hundreds of events at a time, while the latter typically needs many thousands of events to be efficient: in this context, achieving higher throughputs with smaller grids would provide one way to use more manageable, but still reasonably efficient, event generation jobs.
More in detail, in this paper I describe my work on four sets of kernel splitting developments: (1) helicity streams; (2) color sum splitting into separate GPU kernels; (3) color sum host refactoring with optional BLAS offloading; and (4) Feynman diagram splitting into separate GPU kernels. The performance of each set of developments was tested for different physics processes, software configurations and hardware architectures. While these developments are almost exclusively aimed at improving throughputs on GPUs, the large code refactoring that they imply was also propagated to the vectorised C++ code for SIMD CPUs, as code generated by the CUDACPP plugin is strictly the same for CPUs and and GPUs, and the distinction between the two cases is only done at build time through #ifdef directives. I therefore analysed not only CUDA and HIP performance on NVidia and AMD GPUs, but also C++ performance on vector CPUs using different SIMD configurations. Considering the results of all my tests, my recommendation is to include the first three enhancements in a new production release of CUDACPP, while keeping BLAS offloading as an option that is disabled by default. The fourth development, conversely, results in a significant performance degradation on GPUs, and to a lesser extent also on CPUs, and my recommendation is to keep this as a proof-of-concept implementation in a separate development branch, in view of possible further refinements of this approach.
As these new developments imply a significant refactoring of the existing CUDACPP code base, and rely on internal features that have not yet been documented, I also found it useful to start this article with a brief review of some architectural aspects of the software. In this context, I should mention that the internal design of the CUDACPP plugin, which was largely my own work, is far from perfect and the code could certainly benefit from an extensive cleanup. In particular, over time I have added various switches to support alternative parallel implementations of some software components: I believe that this has been very useful, and in some cases essential, to try out different approaches and eventually converge on well-defined solutions to address specific problems, but I am aware that this has resulted in code that might be more difficult to read, and where some of this flexibility may now represent an unnecessary complication. I hope that this article may help clarify the rationale behind some design choices and provide guidance for possible future developments and/or cleanups. I stress, in any case, that the opinions that I express in this paper are only my own, and may differ from those of some of my colleagues.
The outline of this paper is the following. In Sec. 2, I briefly review some aspects of the CUDACPP software architecture design, focusing on those that are most relevant to kernel splitting. In Sec. 3, I describe in detail the software design and implementation of the various kernel splitting enhancements, and I provide the results and an analysis of various performance tests. In Sec. 4, I give my conclusions and an an outlook for this work.
Figure 1: Schematic representation of the architectural evolution of the work on the MG5aMC CUDACPP plugin between 2020 and 2024. This plot was prepared for CHEP2024 and is taken as-is from its proceedings bib:chep2024 , where additional details can be found.
2 Software architecture review
The overall software architecture design of the CUDACPP plugin for GPUs and vector CPUs, its evolution over time, its interplay with the MG5aMC framework and many specific implementation choices have already been described in the previous papers bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 from our team and I will not repeat this here. The main point of this section, instead, is to give further information about some software aspects which have not yet been documented and which are relevant to kernel splitting, such as memory layouts and memory access. First, in any case, I will give a very high level overview of the various areas and phases of development on CUDACPP, to put the kernel splitting work into context.
2.1 Three software areas and development phases
The evolution of the work on the CUDACPP plugin between 2020 and 2024 is represented schematically in Fig. 1 and was summarised in detail in the CHEP2024 proceedings bib:chep2024 , from which that plot is taken. Many more details about the progress of this work over time, including snapshots and links to the initial presentations of some of these ideas at working group meetings of the Madgraph on GPU project, can be found in the slides of my CHEP2024 presentation bib:chep2024slides .
The main aim of the CUDACPP software, from the very beginning, has been to use event-level data parallelism to try and speed up the calculation of MEs, which is the computational bottleneck of the whole MG5aMC framework. This is because, to a large extent, the same mathematical functions need to be numerically computed for different events during the ME calculation, and this makes it possible to efficiently compute the MEs for many events in lockstep at the same time using GPUs and SIMD CPUs. This is something that I had pointed out in two presentations to the HSF generator working group bib:hsf2020 and the LHCC bib:lhcc2020 in 2020, using plots from which I later derived those in Fig. 1.
Largely speaking, the work on the CUDACPP plugin has concerned three areas of the software, roughly corresponding to three phases of development:
  1. 1.
    ME engines: the port to GPUs and SIMD CPUs, using CUDA/HIP and vectorised C++, of the computational engines to calculate MEs from particle momenta, and their backport to the code-generating Python framework, for different physics processes.
  2. 2.
    MadEvent integration: the injection of the new code into one madevent application, by modifying the existing Fortran and linking the new ME engine as a C++ library into it.
  3. 3.
    Full workflow orchestration: the integration and testing of the full MG5aMC workflow involving many madevent applications, including the packaging and installation of the CUDACPP plugin.
Over time, the focus of developments has gradually shifted from the first to the last of these software areas, even if some work on ME engines and on the integration of C++ and Fortran continued up until the release in October 2024. This is reflected in the documentation provided in our previous papers. What I want to stress here, to put it into context, is that the new kernel splitting work that I describe in Sec. 3 was all done at the level of the ME engines alone, even if its motivation (as is the case for all of the work in the Madgraph on GPU project) lies in the streamlining of the full MG5aMC workflow. I mention this because I believe that, amongst the many previous documents from this project, those describing the software features most relevant to kernel splitting in the ME engines are actually the very first ones, namely the proceedings bib:vchep2021 and presentation bib:vchep2021video from the vCHEP2021 conference.
More specifically, in all of my work on kernel splitting presented in Sec. 3, I only developed code using the standalone application check.exe, where all three main components of an event generation application (random numbers, phase space sampling and ME calculation) are implemented in C++ and CUDA using a data parallel approach. This is because check.exe makes it possible to focus on the optimization of the ME engine without being constrained by the MadEvent Fortran infrastructure, where the legacy phase space sampling and other sequential non-ME components slow down the whole workflow. I routinely built check.exe using the CUDACPP makefile cudacpp.mk, without even building MadEvent or any other Fortran code (which proceeds through a separate makefile that internally delegates the build of the CUDACPP library to cudacpp.mk).
int main(…) // application check.exe (similar code in class Bridge)
int nevt = m_gpuBlocks * m_gpuThreads; // number of events == GPU grid size
DeviceBufferMomenta m_devMomenta( nevt ); // memory buffer for nevt events
DeviceBufferMatrixElements m_devMEs( nevt ); // memory buffer for nevt events
// Compute MEs for nevt events (m_gpuBlocks * m_gpuThreads) in one go
MatrixElementKernelDevice mek( m_devMomenta, m_devMEs, m_gpublocks, m_gputhreads, );
mek.computeMatrixElements( );
void MatrixElementKernelDevice::computeMatrixElements( )
gpuLaunchKernel( sigmaKin, m_gpuBlocks, m_gpuThreads, m_momenta.data(), m_MEs.data(), );
__global__ void
sigmaKin( const fptype* momenta, // input: momenta[nevt*npar*4]
fptype* MEs, ) // output: MEs[nevt], final sum over ihel
const int ievt = blockDim.x * blockIdx.x + threadIdx.x; // one event <-> one thread
// Zero the running sum MEs[ievt] and add the contribution of each helicity ihel
MEs[ievt] = 0;
for( int ihel = )
calculate_wavefunctions( ihel, momenta, MEs, );
// Randomly select a color and a helicity for event ievt
selcol[ievt] = …; selhel[ievt] = // extra output (from extra input random numbers)
__device__ void
calculate_wavefunctions( int ihel,
const fptype* momenta, // input: momenta[nevt*npar*4]
fptype* MEs, ) // output: MEs[nevt], running sum over ihel
using M_ACCESS = DeviceAccessMomenta; // non-trivial access (nevt events)
using E_ACCESS = DeviceAccessMatrixElements; // non-trivial access (nevt events)
using W_ACCESS = DeviceAccessWavefunctions; // trivial access (one event)
using A_ACCESS = DeviceAccessAmplitudes; // trivial access (one event)
fptype* wf[nwf] = …; // local variable for one event (wavefunctions)
cxtype amp[1]; // local variable for one event (amplitude for one Feynman diagram)
cxtype jamp[ncolor] = {}; // local variable for one event (dual amplitudes for color flows)
// Feynman diagram 1 of 2
vxxxxx<M_ACCESS, W_ACCESS>( momenta, ihel, …, wf[0] ); // compute wf[0]
vxxxxx<M_ACCESS, W_ACCESS>( momenta, ihel, …, wf[1] ); // compute wf[1]
oxxxxx<M_ACCESS, W_ACCESS>( momenta, ihel, …, wf[2] ); // compute wf[2]
ixxxxx<M_ACCESS, W_ACCESS>( momenta, ihel, …, wf[3] ); // compute wf[3]
VVV1P0_1<W_ACCESS, …>( wf[0], wf[1], …, wf[4] ); // compute wf[4]
FFV1_0<W_ACCESS, A_ACCESS, …>( wf[3], wf[2], wf[4], &amp[0] ); // compute amp[0]
jamp[…] += * amp[0];
// Feynman diagram N of N
// compute wavefunctions and amplitudes…
// Add ME contribution for helicity ihel (from quadratic form on color amplitudes)
fptype& ME = E_ACCESS::kernelAccess( MEs ); // ME for event ievt (from threadIdx.x etc)
ME += sum_ij ( cxconj((cxtype2)(jamp[i])) * colormatrix[i][j] * (cxtype2)(jamp[j]) );
Table 1: Pseudo-code of the ME calculation in the latest CUDACPP, before any kernel splitting changes. Function main represents the check.exe standalone application. Note that sigmaKin is the only GPU kernel (__global__ function) in this workflow. Only a simplified GPU version of the code is shown, using scalar fptype floating point and cxtype complex number types: the actual code that supports both GPUs and SIMD CPUs uses fptype_sv and cxtype_sv data types inside calculate_wavefunction, making it possible to use the same formal code for scalar data and for SIMD vectors (see Ref. bib:vchep2021 for details). The CPU branch of the code for main, encapsulated by an #ifdef and not shown here, uses the Host versions of various classes instead of the Device versions. The pseudo-code for the color sum schematically indicates that this proceeds via a quadratic form using a color matrix, and that color amplitudes are converted from fptype/cxtype floating point precision to a potentially different fptype2/cxtype2 precision (this enables the “mixed” precision mode, where the former is based on double and the latter on float).
2.2 A closer look at the internals of the ME engine
Consider a physics process where npar is the total number of initial and final state particles. Assume that the calculation of Feynman diagram amplitudes is performed using floating types fptype, which is a typedef for either double or float. Very schematically, as shown in Fig. 1, the ME computational engine of CUDACPP is simply a software component that takes as input an array of particle momenta for many events (i.e. 4*npar fptype values for each of nevt events), and returns as its output an array of matrix elements (i.e. one fptype value for each of nevt events). The calculation in CUDACPP is actually more complex than this, as the ME engine also receives other inputs, such as: an event-by-event running coupling $\alpha_{s}$ (at a scale which is computed from particle momenta using an external Fortran module); two event-by-event random numbers for the selection of one specific helicity and leading color in the generated event (which are also returned as outputs); and an event-by-event channel identifier (for MadEvent single-diagram enhancement, also resulting in additional outputs). However, I will neglect this in the following, except otherwise stated, and focus only on the calculation of output MEs from input momenta.
In the current production version of the CUDACPP plugin, v1.00.02 in MG5aMC v3.6.3, the ME computational engine on GPUs is essentially a single and very large monolithic GPU kernel, named sigmaKin. This performs all of the relevant operations that are needed to compute the output ME from the input momenta for each event (one notable exception being the calculation of other couplings that depend on $\alpha_{s}$, which is delegated to a separate kernel). Very schematically, this is represented in the pseudo-code in Table 1. In particular, sigmaKin performs the following operations for each event: it keeps a running sum of the ME over helicities; it loops on all possible helicities of the external particles (or, more precisely, only on a pre-determined set of “good” helicities whose contribution is non-zero); it adds to the running sum of the ME the contribution from each helicity, by calling a device function calculate_wavefunctions which internally involves for each QCD color flow the calculation of particle and propagator wavefunctions and that of dual amplitudes, followed by the sum over all color flows using a quadratic form based on a color matrix; after the end of the helicity loop, sigmaKin randomly draws a helicity and a color for the generated event based on the ME contributions from individual helicities and colors. The calculation of helicity amplitudes for Feynman diagrams, in particular, is described in detail in Ref. bib:vchep2021 (see Fig. 1 therein); complementary details are also given in Ref. bib:zw2025 . As shown schematically in Table 1, this involves three types of elements to compute: the wavefunctions of external (initial/final state) particles, via functions like VXXXXX; the wavefunctions of internal propagators, e.g. via VVV1P0_1; the dual amplitudes for a given colour flow, e.g. via FFV1_0. The names and roles of these functions in CUDACPP are the same as in the original HELAS bib:helas and ALOHA bib:aloha implementations.
One key aspect in the design of this software chain, which has not been documented in detail so far, is the allocation of memory buffers and their access and use in computational kernels, in both the GPU and SIMD CPU code. This was a core part of my design and implementation work in the second half of 2020, which was motivated mainly by the aim of achieving SIMD speedups on CPUs, but also by the aim of achieving coalesced memory access on GPUs, and turned out to be a key enabling factor for both. The design principle which I adopted, in particular, is a complete separation of three key elements of the software (for details, see the backup slides 70–75 from December 2020 in Ref. bib:chep2024slides ):
  1. 1.
    Memory allocation. In the check.exe standalone application (or in the Bridge component that gets linked with madevent), CUDACPP allocates host and device memory buffers that are properly dimensioned for the number nevt of events that will be processed in parallel in one offloading cycle. On the GPU, nevt is simply the GPU grid size where kernels are launched (i.e. the product of the number of threads per block and of the number of blocks), as CUDACPP kernels use event-level data parallelism where each GPU thread processes one event. The classes responsible to allocate and hold the pointers to the memory buffers have no knowledge of the internal memory layout of the buffers; for instance, DeviceBufferMomenta in Table 1 allocates 4*npar*nevt fptype values, but it does not know the layout of the particle momenta arrays. In particular, as opposed to earlier versions of the code that were using structured memory allocations like std::vector in C++ and cudaMalloc3D in CUDA, from the end of 2020 the code uses unstructured memory allocations of raw buffers, both on the host (via malloc) for the GPU and SIMD CPU code, and on the device (via cudaMalloc) for the GPU code.
  2. 2.
    Memory layout and data access. The specific layout chosen for storing data inside the raw memory buffers allocated in the previous step is encapsulated in a separate set of memory access classes. All these classes have methods named kernelAccess, or some variation of this. On the GPU, these methods take as input the fptype* pointer associated to a raw memory buffer for nevt events, and they return as output a data item for the single event ievt processed by the GPU thread where the code is executed, typically indexed by the identifier of the GPU thread, blockDim.x*blockIdx.x+threadIdx.x. In some cases, these methods have additional parameters: method kernelAccessIp4IparConst in the DeviceAccessMomenta class, for instance, takes two additional parameters ip4 and ipar in order to retrieve a given 4-vector component for a given particle. It is only the memory access classes that, internally, are able to decode the memory layouts of a raw buffer: this is done by interpreting the buffers as one-dimensional C-style array or casting them as multi-dimensional C-style arrays. For most data buffers, a Structure-Of-Array (SOA) or an Array-Of-Structure-Of-Array (AOSOA) layout is chosen, where the data items of a given type for different events are contiguous: as explained in detail in the vCHEP2021 proceedings bib:vchep2021 , this is absolutely essential for SIMD processing on vector CPUs, and it is also beneficial – but not at all a strict requirement – on GPUs to improve performance through coalesced data access. In some cases, such as couplings, the floating point raw arrays include the real and imaginary components of complex data, and it is the responsibility of the data access classes to return outputs with the API of a complex number data type. The case of CPU code, which I will not discuss here in detail as it is less relevant to kernel splitting, includes the additional complication of returning outputs whose data types fptype_sv and cxtype_sv can be scalar (for no-SIMD C++) or SIMD vectors (through compiler vector extensions bib:vchep2021 , hereafter CVEs) of fptype or cxtype data. Concerning the use of memory access classes in other CUDACPP software components, notably the computational methods for helicity amplitudes like VXXXXX, VVV1P0_1 or FFV1_0, I initially chose to implement this as a template parameter: in Table 1, for instance, DeviceAccessMomenta is a template parameter M_ACCESS of VXXXXX. The idea behind the choice of using templates was to allow the flexibility of easily switching between different memory layouts at build time, to compare their performances; in retrospective, the same flexibility may have been achieved with the same memory access classes but without templates. One could now consider removing those templates, although it is not clear to me if this would make the code easier to read and maintain, or faster to build.
  3. 3.
    Arithmetic operations and other computational functions. Finally, the decoupling of data access from actual calculations has probably been one of the most important aspects of the whole CUDACPP software design. Just like the memory allocation classes, also the helicity amplitude functions like VXXXXX, VVV1P0_1 or FFV1_0 have almost [^1] One notable exception is that some assumptions are still made about the memory layout of wavefunctions, which for spin 1/2 and spin 1 particles are complex 6-dimensional arrays as in the original HELAS bib:helas and ALOHA bib:aloha implementations, from which CUDACPP is derived. This detail is relevant to the kernel splitting for different Feynman diagrams, as discussed in Sec. 3. no knowledge of the memory layouts used in the multi-event data buffers. These functions also ignore whether the calculations are performed in single or double precision, as I encapsulated this choice in the fptype type definition. Even more, these functions ignore whether the arithmetic operations within them are applied to scalar values on GPUs or CPUs or to SIMD vectors on CPUs: in the latter case, this is possible because simple operators like “+” are automatically understood by the compiler as vector operations when applied to CVE vectors of fptype, and are also implemented using SIMD CVE operations in their definition for vectors of custom types, notably the vectors of cxtype complex types. This design is extremely powerful because it has made it possible to use formally the same exact lines of software in the code-generated helicity amplitude functions like VVV1P0_1 or FFV1_0. I note in passing, however, that the functions computing wavefunctions for initial and final state particles, like IXXXXX, OXXXXX and VXXXXX, or their variations for massless or beam-collinear particles, are not code-generated from a model Lagrangian, but are instead hardcoded and often required particular care and successive iterations. Two complications that I had to address for SIMD code, in particular, are the following [^2] In this context, I find it important to stress that, in my experience, achieving a robust and performant CUDACPP implementation for SIMD CPUs was far from a trivial task, and was, in fact, much more complex than achieving it for GPUs. The difficulties for implementing IXXXXX-like functions are just two examples, and I already mentioned the fact that there are very strict constraints in the memory layouts for CPU SIMD, unlike those for GPUs. Another example is the implementation of mixed floating point mode, where I added a complex mechanism to merge/split two SIMD vectors of double into/from one SIMD vector of float at the boundary between dual amplitude and color sum calculations, in order to use the widest possible SIMD vectors in both cases and maximize efficiency. More generally, in my opinion, it is fair to say that achieving SIMD speedups through vectorization is always much more complicated than achieving speedups on GPUs. Code acceleration on GPUs, in fact, can be achieved even without lockstep processing (at the cost of some inefficiency from warp divergence), or without optimized memory layouts (at the cost of some inefficiency from the lack of coalesced data access), simply because a GPU has thousands of threads. One good feature of CUDACPP is that, by focusing on achieving C++ vectorization on SIMD CPUs, this has also improved the efficiency of the code on GPUs. : first, unlike VVV1P0_1 or FFV1_0, the IXXXXX-like functions include some if/else branching, which I reimplemented using vector masks in the SIMD case in order to fully exploit data parallel speedups (even if, from a performance point of view, this is only important for simple processes with few Feynman diagrams); second, as I performed all my functional testing with Floating Point Exception traps (FPEs) enabled in order to develop more robust code, I came across some crashes caused by the interplay of compiler optimizations, SIMD CVEs and FPE traps, which I addressed using volatile keywords and which pushed me to develop a large set of functional tests specifically for these functions.
On top of the separation of the three software concerns above (data allocation, data access and arithmetic calculations), another important part of the design of the CUDACPP software was the implementation of kernel launching on GPUs and event loops on CPUs. My aim here was to keep as much as possible of the software logic and of the actual code identical for GPUs and CPUs, in order to simplify the iterative addition of new features and of bug fixes for both cases. In short, I addressed this by keeping also for SIMD CPUs the idea of a “grid” of events that are processed in one given iteration, and by explicitly subdividing this grid into SIMD event vectors and adding an explicit loop over them. Much more practically, and more importantly for the kernel splitting work described in the next section, the link between memory allocations and the execution of the ME computational engines is provided by two different incarnations of a MatrixElementKernel (MEK) class, one for GPUs (Device) and one for CPUs (Host). A singleton instance of the appropriate MEK class is constructed and used inside check.exe for standalone tests, or inside the Bridge component that is linked with Fortran in madevent for the full MG5aMC workflow. As shown in Table 1, it is the MatrixElementKernelDevice class that internally launches the monolithic sigmaKin kernel in the current CUDACPP. This is the starting point for the enhancements presented in the following section.
Process Feynman diagrams SDE channels Leading colors Color matrix Helicities Wave functions LOC (CPPProcess.cc)
$gg\!\rightarrow\!t\bar{t}$ 3 3 2 2 x 2 16 5 1424
$gg\!\rightarrow\!t\bar{t}g$ 16 15 6 6 x 6 32 12 1641
$gg\!\rightarrow\!t\bar{t}gg$ 123 105 24 24 x 24 64 26 3569
$gg\!\rightarrow\!t\bar{t}ggg$ 1240 945 120 120 x 120 128 121 31140
Table 2: Comparison of the computational complexity of the four physics processes considered in the tests described in this paper. The columns represent the following: (1) the physics process; (2) the number of Feynman diagrams; (3) the number of distinct SDE channels, i.e. the number of diagrams used in the MadEvent single diagram enhancement algorithm; (4) the number of leading QCD color flows; (5) the size of the color matrix; (6) the number of combinations of helicities for initial and final state particles (for the four processes considered, the number of good helicities contributing to the ME calculation coincides with the total number of helicities); (7) the number of wavefunctions to be computed for all external (initial and final state) and internal (propagator) particles; (8) the number of lines of code in the CUDAPP generated file CPPProcess.cc, in the ihel3 version of the software. All of these numbers are needed to describe the complexity of the ME calculation in a single madevent (or standalone) application, which is the main focus of this paper; the number of SDE channels, in addition, is one of the factors that determines how many instances of a madevent application must be launched in a full event generation workflow. The number of helicities determines how many GPU streams are used (in the ihel1 software). The number of leading colors determines the number of dual amplitudes that are computed using Feynman diagrams and used as input to the color sum; as of the ihel2 software, these are stored into and retrieved from GPU global memory. The size of the color matrix determines the complexity of the color sum calculation (addressed with a separate GPU kernel in the ihel2 software, and optionally via BLAS in the ihel3 software). The number of wavefunctions is particularly relevant to the ihel4 software, as these are stored into and retrieved from GPU global memory in that case.
3 New developments: GPU kernel splitting
As discussed extensively above, a key parameter of the CUDACPP ME computational engine is the number nevt of events that are processed in parallel in one offloading cycle (in practice: in one call of the MEK computeMatrixElements function in Table 1). On a GPU, this is simply equal to the GPU grid size, i.e. to the product of the number of blocks gpuBlocks and of the number of threads per block gpuThreads. All relevant multi-event arrays, both in the Fortran madevent application and in the CUDACPP Bridge component (or in the check.exe standalone application), must be large enough to contain nevt events: the larger the GPU grid size, in particular, the larger the RAM footprint of the application (which can become very large bib:ichep2022 for madevent). This parameter can be configured at build time (and partly at runtime) in the madevent application, where it is referred to as VECSIZE bib:zw2025 , and is highly configurable in check.exe, where gpuBlocks and gpuThreads are defined independently at runtime.
As I briefly hinted above, one of the main aims of the new developments presented in this section is the fact that the ME calculation throughput (in MEs/s) on a GPU is generally quite low for small grids i.e. small values of nevt, and only reaches a peak plateau for relatively large grids, i.e. large values of nevt. This is an issue that exists in CUDACPP since the very beginning of our developments: in Fig. 5 of our vCHEP2021 proceedings bib:vchep2021 ), for instance, we had shown that the ME throughput for the $gg\!\rightarrow\!t\bar{t}gg$ process only ramps up significantly with at least 16k events in the grid, and reaches the peak plateau even later, with 128k events in the grid. In the same paper, we had also already suggested that higher throughputs, possibly because of lower “register pressure”, might be achieved by splitting monolithic sigmaKin kernel into smaller kernels: two specific ideas that we had mentioned, in particular, were the possible use of different GPU threads to process different helicities in the same event, or the possible use of CUDA Graphs to orchestrate a much larger number of smaller GPU kernels. I will come back to both of these ideas in the following.
Since the ramp-up of ME throughputs with increasing GPU grid sizes is one of the main issues that this new work aims at addressing, or in any case a very good test of the effectiveness of these code changes, the results that I present in this section will essentially all consist of plots of that sort. One limitation of this work is that, due to lack of time, I will only show plots for a single CPU process, and only for the standalone application check.exe. This is unfortunate, because the practical benefits of any progress in this area would mainly be for the full MG5aMC workflow, using several madevent applications accessing the GPU in parallel. In the ACAT2022 presentation bib:acat2022 (see Fig. 1 therein), I had also studied the ramp-up of throughput as a function of the GPU grid size when several check.exe applications are launched in parallel, which seemed to indicate that the use of the GPU is more efficient in that case. Those results had been obtained using software containers prepared for the HEP-SCORE benchmarking project bib:bmk ; similar tests may be repeated in the future, when updated HEP-SCORE containers are preppared using more recent versions of CUDACPP, possibly including those presented in this paper.
Concerning my software development process, I note that for this new research I used the same methodology that I have been following for all of my work on CUDACPP during the last six years. In particular, I always prototyped, tested, fixed and optimized my changes using as baseline the code-generated software for one specific physics process, typically $gg\!\rightarrow\!t\bar{t}$ or $gg\!\rightarrow\!t\bar{t}gg$. In the case of this work on kernel splitting, I focused on the CUDA implementation but also tested the SIMD C++ version. I then committed all changes for the given process to git, until these reached a state that I considered reasonably complete. At that point, I backported my changes to the code-generating Python framework, and regenerated all physics processes. I then performed larger-scale functional and performance tests, also using different hardware implementations like AMD GPUs, and iterated until completion. This development process has been possible only because, while the git repository of the MG5aMC project bib:mg5amc-github mainly contains the hardcoded components of the framework and its code-generating engine, the git repository of madgraph4gpu project bib:mg4gpu-github contains not only the CUDACPP plugin but also the code-generated software for several physics processes. In my experience, this has been a key ingredient in the development of these latest enhancements to CUDACPP, but also much more generally of the whole plugin. In particular, I stress that I would never have been able to design and implement such large code changes directly in the code-generating Python code.
Figure 2: Schematic representation of the CUDACPP engine for computing MEs (sigmaKin), and of its evolution through the first four scenarios described in this paper: (ihel0) current version before kernel splitting; (ihel1) helicity streams; (ihel2) color sum kernel; (ihel3b) color sum on BLAS via host dispatcher. For the ihel3 software, only the (non-default) case with BLAS enabled at runtime is illustrated: by default, the ihel3 software has BLAS disabled at runtime, which is essentially the same as what is shown for the ihel2 scenario (the only difference is that in the ihel3 scenario the kernel is named color_sum_kernel and is invoked by a color_sum_gpu host function, which could also dispatch the calculation to the color_sum_blas BLAS host function).
Figure 3: Schematic representation of the CUDACPP ME engine (sigmaKin), in the last kernel splitting scenario described in this paper: (ihel4) Feynman diagrams as individual kernels.
Figure 4: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU (node itscrd90 at CERN). Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. The number of threads per block is fixed to 32 (NVidia GPU warp size); the grid size is varied by changing the number of blocks. Each plot compares the different scenarios considered in this paper: (ihel0) current version before kernel splitting; (ihel1) helicity streams; (ihel2) color sum kernel; (ihel3) color sum kernel via host dispatcher; (ihel3b) color sum on cuBLAS, without TF32 math mode; (ihel4) Feynman diagrams as individual kernels.
Figure 5: Throughputs (ME/s) as a function of grid size for an AMD MI200 GPU (at LUMI). Higher is better. The 9 plots correspond to 3 physics processes in 3 floating point precisions. The number of threads per block is fixed to 256; the grid size is varied by changing the number of blocks. Each plot compares the different scenarios considered in this paper: (ihel0) current version before kernel splitting; (ihel1) helicity streams; (ihel2) color sum kernel; (ihel3) color sum kernel via host dispatcher; (ihel3b) color sum on hipBLAS via host dispatcher; (ihel4) Feynman diagrams as individual kernels.
In the following subsections, I will describe the four sets of kernel splitting changes I developed: (ihel1) helicity streams; (ihel2) color sum splitting into separate GPU kernels; (ihel3) color sum host refactoring with optional BLAS offloading; and (ihel4) Feynman diagram splitting into separate GPU kernels. These developments are sequential, i.e. ihel4 includes ihel3, which includes ihel2, which includes ihel1. The software architecture of the sigmaKin ME engine is represented in Fig. 2 for the first three scenarios (where it is also compared to the situation before any kernel splitting, hereafter “ihel0”) and in Fig. 3 for the ihel4 case. Note that I will not provide new pseudo-code listings for my four kernel splitting developments, but in some cases I will refer to the pseudo-code for the current version ihel0 in Table 1 to point out what I changed or which technical issues I had to address. The results of my tests for these different versions of the software, which I will discuss more in detail in each subsection, are given in Fig. 4 for an NVidia V100 GPU at CERN and in Fig. 5 for an AMD MI200 GPU at LUMI, for three floating point precisions (double, mixed, float) and for four physics processes of increasing complexity ($g\!\rightarrow\!t\bar{t}, gg$$\rightarrow\!t\bar{t}g, gg$$\rightarrow\!t\bar{t}gg, gg$$g\!\rightarrow\!t\bar{t}ggg$), whose relevant parameters are described in detail in Table 2.
3.1 Helicity streams (“ihel1”)
The internal substructure of the sigmaKin ME engine in the current version (ihel0) of CUDACPP is illustrated schematically in the top-left diagram in Fig. 2. This gives a visual representation of the pseudo-code in Table 1. As mentioned previously and as visible in the diagram, the two main components of the calculation for each event, namely the computation of wavefunctions and dual amplitudes for a given color flow from Feynman diagrams, and their squaring and sum over all color flows, are performed sequentially for all helicities. This workflow makes it impossible to split these two components into separate kernels, because the loop over helicities is effectively inside the loop over events.
The very first step in my kernel splitting developments was therefore, quite naturally, to reverse this situation and make the event loop the innermost loop, inside an outermost loop on helicities. This was the focus of my work with Olivier Mattelaer during the 2022 GPU hackathon (and is quite likely an idea that he suggested, in fact). In practice, the main change to achieve here was to turn sigmaKin from a __global__ device kernel into a host function, and to turn instead calculate_wavefunction from a __device__ function callable by a kernel into a kernel itself. Other computations also had to be modified: for instance, the selections of event-by-event colors and helicities, which the sigmaKin kernel was performing outside the helicity loop (see Table 1) have now also been turned into separate GPU kernels. These changes, which I achieved in November 2024 bib:ks2024 , by themselves already provide a moderate increase of throughputs for small grids.
The real breakthrough, however, came when the parallel calculations for different helicities were moved to separate CUDA Streams. In this “ihel1” version of the software, which is illustrated in the top-right diagram in Fig. 2, ME throughputs on NVidia GPUs reach their peak performance with much smaller grids than in the current ihel0 version, for all physics processes and floating precisions I tested. The peak throughput themselves are also increased by around 10-20%. This can be seen in Fig. 4, by comparing the blue (ihel0) and orange (ihel1) curves. The improvement is especially impressive for complex processes like $gg\!\rightarrow\!t\bar{t}ggg$, where peak throughputs are reached with $\cal O$}(100) events per grid in ihel1, as opposed to $\cal O$}(10k) in the current ihel0. These improvements from ihel1 are essentially the single most important progress described in this paper. Most likely, the improvement comes from the increase in parallelism of the workflow: instead of a single kernel launch in each offloading cycle, which runs for a long time because it internally loops over helicities, there are now several, much shorter kernels launched in parallel, one for each helicity, in a separate GPU stream for each helicity.
For AMD GPUs, where the same solution was implemented using HIP Streams, the benefits are much less clear, as shown in Fig. 5: throughputs increase with both small and large grids for complex processes like $gg\!\rightarrow\!t\bar{t}gg$, but for simpler processes the opposite effect is observed. The code refactoring in ihel1 was also propagated to the SIMD C++ code, and tested on a reference Intel node, which has excellent AVX512 performance thanks to the presence of 2 FMA units. I will not give detailed numbers, but my tests show that the throughputs are essentially unchanged for all SIMD builds, within the expected 1-2% fluctuations.
3.2 Color sum as a separate GPU kernel (“ihel2”)
The next step in my kernel splitting developments consisted, quite naturally, in separating the two main components of the ME calculation, namely, (1) the calculation from Feynman diagrams of the dual amplitudes for ncolor color flows (jamp: a vector $J$ with ncolor complex elements) and (2) the color sum. Initially, I simply split the calculate_wavefunction kernel, which was doing both computations in ihel1, into two separate kernels: calculate_jamps, which calculates the jamp $J$, and color_sum, which computes the quadratic form $J^{H}(C)J$ for a symmetric real color matrix $C$ using the vector $J$ and its conjugate transpose $J^{H}$. This “ihel2” version of the software is illustrated in the bottom-left diagram in Fig. 2. One important difference in this case is that the jamp variable is no longer a local variable inside calculate_wavefunction, as it was in ihel0 (see Table 1) and ihel1: instead, it is now a GPU global memory buffer that is allocated outside the MEK component and is accessed, with the appropriate layout decoding provided by a new class DeviceAccessJamps, in both the calculate_jamps and color_sum kernels.
As can be seen in Fig. 4, by comparing the orange (ihel1) and green (ihel2) curves, on Nvidia GPUs this further change in the software yields an additional increase of the peak throughput by 10-20% for complex processes like $gg\!\rightarrow\!t\bar{t}gg$ and $gg\!\rightarrow\!t\bar{t}ggg$. For simpler processes, conversely, it results in a minor decrease of peak throughputs; in my opinion, this is a moderate cost that can be tolerated, as speeding up complex processes is more important. On AMD GPUs, as can be seen in Fig. 5, the additional change results in a minor increase in peak throughputs for $gg\!\rightarrow\!t\bar{t}gg$, but for simpler processes ihel2 is almost indistinguishable from ihel1. In the CPU implementation, throughputs are again essentially unchanged for all SIMD builds within 1-2%.
3.3 Color sum dispatcher to kernel or BLAS (“ihel3”)
The next step of my developments consisted in investigating the possible use of the cuBLAS linear algebra library for computing color sums on GPUs, instead of using a CUDA kernel. One of the main motivations for this work was that the current CUDACPP code only uses the traditional CUDA Cores, but a large part of the computing power on recent NVidia GPUs comes from specialized Tensor Cores designed for the matrix algebra operations used in AI, and cuBLAS may provide a way for CUDACPP to exploit them. Developing code for Tensor Cores, in fact, is challenging because it requires the use of programming APIs other than CUDA: an easier alternative consists in using specialized libraries for AI or linear algebra that internally use the Tensor Core APIs, cuBLAS being one of them.
The BLAS implementation I developed is fully integrated in CUDACPP, and its functionality has been extensively tested both in the standalone use case and in the full MG5aMC workflows. It was originally developed for cuBLAS on Nvidia GPUs, but it has also been ported to AMD GPUs using the hipBLAS wrapper for rocBLAS. Studies of standalone color sums with cuBLAS and Tensor Cores had already been done bib:srblas in previous years by my colleagues in the Madgraph on GPU project: the work that I present in this section, however, is not based on the code developed for those studies and represents a restart from first principles.
In practice, my work on this “ihel3” version of the software was the following. To start with, I encapsulated the color sum calculation on GPUs in a host function color_sum_gpu (in parallel, I also created a color_sum_cpu function for the vectorized C++ version on SIMD CPUs). To make the software more modular and more manageable, I also took this opportunity to clean up the color sum code and move it to a separate source code file. The color_sum_gpu host function is just a wrapper that may dispatch the color sum calculation to two different GPU implementations:
  1. 1.
    color_sum_kernel. By default, the calculation is performed using a kernel color_sum_kernel. This is essentially the same code as in the ihel2 software, with one minor difference: the GPU global memory layout of the dual amplitudes, which was an SOA jamp[ncolor][2][nevt] in ihel2, is now an SOA jamp[2][ncolor][nevt], because this makes it easier to separate the real and imaginary parts of the dual amplitudes for the BLAS calculation, and the same layout is used for simplicity in both the kernel and BLAS implementations. As can be seen in Fig. 4 by comparing the green (ihel2) and red (ihel3) curves, the achieved performance is indistinguishable from that of ihel2.
  2. 2.
    color_sum_blas. The second implementation of the color sum consists in a host function color_sum_blas that internally calls the BLAS library. Specifically, since the color matrix $C$ is real and symmetric, the color sum $J^{H}(C)J$ over the vector of dual amplitudes $J=A+iB$ may be decomposed as
    $(A^{t}-iB^{t})(C)(A+iB)=A^{t}(C)A+B^{t}(C)B,$ (1)
    i.e. as the sum of two quadratic forms $V^{t}(C)V$, where the real vector $V$ may represent either the real part $A$ or the imaginary part $B$ of the complex dual amplitude vector $J$. Each calculation involves two steps. In mixed or float precision, for instance, the vector $(C)V$ is computed using cublasSgemm, while its dot product with $V^{t}$ is then computed using cublasSgemmStridedBatched. These two functions, as well as their double-precision and their HIP counterparts, are called through their abstractions via #define directives, in the header-only approach described in Ref. bib:chep2024 . The intermediate results of the first calculation, $(C)V$, are stored in GPU global memory using an additional buffer allocated in the MEK component (the allocation is done at runtime after determining the number of “good” helicities in each physics process, since the amount of memory allocated is proportional to the number of good helicities). As shown in the bottom-right diagram in Fig. 2, the BLAS calculations for different helicities are performed in separate GPU streams: technically, many BLAS handles are used, each associated to a different stream. Since the performance of the BLAS implementation of the color sum is generally worse than that using kernels, as discussed below, this is only available as an option, which must be explicitly enabled at runtime by setting an environment variable [^3] Set environment variable CUDACPP_RUNTIME_BLASCOLORSUM at runtime to use the cuBLAS (on Nvidia GPUs) or hipBLAS (on AMD GPUs) implementation of color sums.. Finally, as the BLAS library contains several switches targeting tensor cores, I also added another environment variable [^4] Set environment variable CUDACPP_RUNTIME_CUBLASTF32TENSOR at runtime to encourage cuBLAS to use Tensor Cores in color sums. This sets math mode CUBLAS_TF32_TENSOR_OP_MATH. to encourage BLAS to use Tensor Cores in the color sum (TF32 math mode).
The performances of the cuBLAS and kernel implementation of color sums on NVidia GPUs are compared in Fig. 4, where they are referred to as “ihel3b” (purple) and “ihel3” (red), even if they both result from the same ihel3 code, with and without an environment variable set at runtime. The picture clearly shows that BLAS performs much worse than CUDA kernels for the simpler $g\!\rightarrow\!t\bar{t}, gg$$g\!\rightarrow\!t\bar{t}g$ and $gg\!\rightarrow\!t\bar{t}gg$ physics processes. For the more complex $gg\!\rightarrow\!t\bar{t}ggg$ process, the situation is more complex: for small grids, the kernel implementation is faster for all floating point precisions; for large grids, however, the BLAS implementation eventually becomes as fast as the kernel implementation, and in double and float precision (but not in mixed precision) it is eventually faster for very large grids. This is interesting, but of not much practical relevance, as most production workflows would use small grids to keep event generation jobs more manageable. On AMD GPUs, Fig. 5 shows that the kernel version of color sums is always much better than the corresponding hipBLAS implementation. Keeping all these observations into account, it seems appropriate to compute color sums using kernels instead of BLAS by default.
Figure 6: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine, for different physics processes and software configurations, in the ihel3 version of the software. The 3 plots correspond to 3 floating point precisions. Measurements performed on a node equipped with an NVidia V100 GPU and an Intel Xeon Silver 4216 CPU (which supports AVX512 but gives better results for CUDACPP in the “512y” mode than in the “512z” mode because it has only one FMA unit). The results shown for CUDA are based on the default ihel3 implementation of color sums using GPU kernels.
Figure 7: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine for the $gg\!\rightarrow\!t\bar{t}ggg$ process, in the CUDA kernel (ihel3) and CUDA/BLAS (ihel3b) implementations, as a function of the number of events nevt processed in parallel. The 3 plots correspond to 3 floating point precisions. For the CUDA kernel implementation, nevt is the product of the number of threads per block, which is fixed to 32, by the varying number of blocks in the GPU grid. For the CUDA/BLAS implementation, nevt is the number of events in the dual amplitude vectors on which the BLAS color matrix multiplication is computed. Measurements performed for an NVidia V100 GPU, on the same node used to prepare Fig. 6.
3.3.1 Color sum profiling
To put this work on color sums into context, and to better understand the relative merit of the cuBLAS and kernel implementations in $gg\!\rightarrow\!t\bar{t}ggg$, I found it useful to perform some more detailed profiling of this calculation. One of my aims was to measure the time taken by the color sum as a fraction of the total time taken by the ME calculation in sigmaKin, in different situations. In fact, the motivation for many recent efforts to speed up the color sum calculation, such as our development of the mixed precision mode bib:acat2022 ; bib:chep2024 ; bib:zw2025 or the work I present here on BLAS, is that the profiling bib:kiran of earlier versions of MG5aMC had shown that this could represent up to 60% of the total ME computation for $gg\!\rightarrow\!t\bar{t}ggg$; however, more recent versions of the software, notably CUDACPP, have not yet been profiled in detail.
Initially, I profiled the code by a sampling approach, using perf, but this did not allow detailed color sum profiling on the GPU. I therefore made some additional modifications to the ihel3 version of the code to instrument it with dedicated timers. Specifically, I used some timers based on the x86 rdtsc instruction, which I had developed for some previous profiling work on MG5aMC bib:prof2024 . This approach can provide relatively accurate results with a limited (<<10%) overhead.
The results of my analysis, which are shown in Fig. 6 for the CUDA and SIMD C++ backends, are somewhat surprising: for $gg\!\rightarrow\!t\bar{t}ggg$, in the ihel3 version of the software, the color sum implementation in CUDACPP represents only 5 to 10% of the total time to compute the ME in all SIMD CPU modes, while in the CUDA kernel implementation this fraction is around 28%, 15% and 6% in double, single and mixed precision modes, respectively. These results are particularly good for the mixed precision mode, as they indicate that our previous optimizations have managed to reduce the time taken by the color sum to a level where this is no longer a bottleneck of the ME calculation. In retrospective, as a consequence, the benefits that one may hope to obtain on GPUs from BLAS are quite limited.
In Fig. 6, the results for BLAS color sums are not shown: this is because they heavily depend on the number of events nevt processed in each GPU cycle, which corresponds to the size of the event vector in BLAS and to the GPU grid size for the calculate_jamp kernels. This is illustrated in Fig. 7, which shows that the efficiency of BLAS color sums increases as nevt increases, while that of the kernel implementation is essentially constant. This plot is consistent with the trend shown in Fig. 4 and is useful to better understand it.
In the C++ implementation, one reason for the relatively low time footprint of the color sum may be a set of optimizations that I introduced in October 2022 and never documented in detail. To start with, as discussed above for BLAS in Eq. 1, I simplified the color sum on the complex vector of dual amplitudes as the sum of two quadratic forms on real vectors. In addition, I also rewrote the multiplication as one involving a triangular matrix, to ensure that the non-diagonal terms of the symmetric matrix are only used once in the calculation.
Color sum implementation Blocks*Threads (nevt) Function FMA Tensor Cores
Kernel 8*1 color_sum_kernel 1.9% 0.0%
Kernel 512*32 color_sum_kernel 9.0% 0.0%
cuBLAS 8*1 ampere_sgemm_32x32_sliced1x4_nt 22.7% 0.0%
gemvNSP_kernel<<…cublasGemvTensorStridedBatched…>> 7.3% 0.0%
cuBLAS 512*32 cutlass::Kernel2<<cutlass_80_simt…>> 81.5% 0.0%
gemvx::kernel<<…cublasGemvTensorStridedBatched…>> 20.6% 0.3%
cuBLAS/TF32 8*1 cutlass::Kernel2<<cutlass_80_tensorop…>> 13.7% 16.4%
gemvNSP_kernel<<…cublasGemvTensorStridedBatched…>> 7.3% 0.0%
cuBLAS/TF32 512*32 cutlass::Kernel2<<cutlass_80_tensorop…>> 14.1% 24.4%
gemvx::kernel<<…cublasGemvTensorStridedBatched…>> 20.3% 0.3%
Table 3: Summary of the results for the profiling of the check.exe application, using the NVidia NSight Compute (ncu) profiler version 2024.1.1.0. The results are obtained on an NVidia A100 GPU, using mixed floating precision in the ihel3 version of CUDACPP. Only the kernels in one specific helicity stream were profiled. The columns represent the following: (1) color sum implementation (three configurations are tested: default using CUDA kernels; cuBLAS; cuBLAS with TF32 math mode); (2) GPU grid size configuration for the kernel implementation (or, equivalently, number of events processed in one cuBLAS multiplication); (3) simplified function name printed out by the profiler; (4) FMA (CUDA Core) activity, as per ncu metric sm__pipe_fma_cycles_active.avg.pct_of_peak_sustained_active; (5) Tensor Core activity, as per ncu metric sm__pipe_tensor_cycles_active.avg.pct_of_peak_sustained_active.
3.3.2 Tensor Core profiling
As the use of Tensor Cores was one of the main motivations for testing the use of cuBLAS for color sums, in Table 3 I present the results of some quick studies using the NVidia NSight Compute (ncu) profiler. Unlike the other results shown in this paper, these were obtained using a more recent NVidia A100 GPU, which has more advanced Tensor Core features than a V100. The results of this test, which was performed using CUDACPP ihel3 with mixed floating precision, clearly show that Tensor Cores are used by cuBLAS color sums. In particular, the cublasSgemm multiplication $(C)V$ makes significant use of Tensor Cores, but only if TF32 math mode is explicitly enabled, as otherwise cuBLAS seems to prefer traditional SIMT kernels. As for the multiplication of $V^{t}$ by $(C)V$ in cublasSgemmStridedBatched, two different implementations are used by cuBLAS for small and large numbers of events, but neither of them makes a significant use of Tensor Cores. Other tests in different configurations, whose results are not presented here in detail, indicate that cuBLAS color sums make a significant use of Tensor Cores (irrespective of whether TF32 math mode is enabled or not) when double precision is used. This is somewhat surprising, as Tensor Cores are designed and optimized for lower precision calculations in AI. More detailed studies could be useful to understand the usefulness of cuBLAS for CUDACPP color sums on more recent NVidia GPUs, with much more advanced Tensor Core features.
3.4 Feynman diagrams as separate kernels (“ihel4”)
The final step in my developments consisted in trying to further decompose the calculate_jamps kernel, which in ihel3 computes dual amplitudes from all Feynman diagrams, into many shorter kernels. The natural way to address this, without major changes in the algorithm, consisted in encapsulating the calculation of every Feynman diagram in an individual GPU kernel. In practice, rather than launching one calculate_jamps kernel per helicity stream, which internally processes all Feynman diagrams, the new code launches many kernels diagram1, diagram2, …diagramN sequentially in each helicity stream. In both cases, the processing of Feynman diagrams must follow a predefined order, because for every diagram the algorithm knows which intermediate wavefunctions have already been computed, and which wavefunctions must still be computed. This “ihel4” version of the software is illustrated in Fig. 3.
Unfortunately, this version of the software has a consistently worse performance than the previous ones, both on NVidia and AMD GPUs, as shown in Figures 4 and 5. With respect to ihel3, for instance, the ihel4 throughputs for $gg\!\rightarrow\!t\bar{t}ggg$ on an NVidia V100 GPU are a factor 3-4 worse than ihel3 at their peak for large grids, and worse by much larger factors for small grids. In addition, ihel4 crashes for very large grids, in configurations where ihel3 performs very well. The ihel4 software in vectorized C++ for SIMD CPUs also performs worse than ihel3, with degradations in throughput of the order of 20% to 50%.
One possible explanation for the worse GPU performance of ihel4 is more frequent memory access. One technical challenge in this case, in fact, is that the wavefunctions for all external and internal particles, which are a local variable wf in the calculate_wavefunctions and calculate_jamps kernels in versions ihel0–ihel3 of the code (see Table 1), must now also be stored into and retrieved from GPU global memory. In addition, as previously mentioned in a footnote, the existing implementation of helicity amplitude functions like FFV1_0 essentially constrains the wavefunction layout in GPU global memory to be an Array-Of-Structures (AOS) wf[nevt][6][2], which is suboptimal as it does not allow coalesced memory access as an SOA would do. This further increases the pressure on the memory bandwidth, which may be one cause of slowdowns.
[…]
One possible way to overcome this issue could be the use of the CUDA Graphs bib:graphsblog technology. This might be a relatively quick option to investigate as a follow-up to this research. Another option could be to group together some Feynman diagrams in kernels of intermediate sizes, but this would probably require larger algorithmic changes. Unless the sequential calculation of Feynman diagrams is reorganised to allow some level of parallelism, in any case, I would not expect these developments to yield throughput increases. Their interest, instead, is mainly that they could make the software more modular and manageable, possibly leading to faster build times and possibly allowing ME calculations with a large number of final state particles. Concerning the calculation of processes with more final state gluons, however, I believe that Berends-Giele recursion relations bib:bg represent a better approach, which has already been successfully ported to GPUs by different generator teams bib:bg-bothmann ; bib:pepper ; bib:martinez .
In summary, taking into account their large performance penalty, I think that the ihel4 developments should only be regarded as an interesting — and fully functional — proof-of-concept, which could possibly represent the basis for further developments. However, they should definitely not be included in a new production release of the CUDACPP software, unlike the ihel1, ihel2 and ihel3 developments, which should instead be merged into the upstream master branch in my opinion.
The software described in this paper is available on github. Two separate tags have been prepared for versions ihel3 bib:ihel3tag and ihel4 bib:ihel4tag . A pull request for the upstream inclusion of ihel3 has also been created.
4 Outlook and conclusions
In this paper, I have described my work on four sets of further GPU optimizations of the CUDACPP plugin for the Madgraph5_aMC@NLO (MG5aMC) generator. The first three optimizations are in my opinion ready for the inclusion in new release of the CUDACPP plugin and of the MG5aMC framework, while the fourth one represents a useful proof-of-concept for further developments. The new developments have targeted NVidia GPUs, but they have also been ported to AMD GPUs and propagated where relevant to the vectorized C++ code for SIMD CPUs. All these changes are fully functional and have been extensively tested on all platforms. I have also taken the opportunity of this paper to describe more in detail some features of the CUDACPP software that are relevant to these new developments and that had not yet been documented.
The new approach mainly consists in splitting the calculation of matrix elements, which had been so far performed using a single large GPU kernel, into several smaller kernels. The first optimization, which parallelizes the calculation for different helicities of the initial and final state particles to different CUDA Streams, achieves a reduction by one to two orders of magnitude in the number of events that must be computed in parallel to make an efficient use of the GPU. This is interesting from a user perspective, as it should allow event generation jobs on GPUs using smaller numbers of events than the current CUDACPP. The second and third optimizations reorganize the calculation of color sums as a separate GPU kernel, which is taken as the default implementation, and possibly as a cuBLAS calculation, which can be enabled at runtime. Delegating the color sum to a separate kernel approximately provides a 10% to 20% improvement in peak throughputs. The cuBLAS implementation has a worse performance in most, but not all, configurations, but it is very interesting as it allows the use of NVidia Tensor Cores, while the existing CUDACPP kernels only use traditional CUDA cores. The fourth development consists in splitting up the computation of different Feynman diagrams in different GPU kernels: its performance is poor, mainly because of kernel launch overheads, but this may represent a useful starting point for further developments, possibly using CUDA Graphs.
Acknowledgements
I thank the MG5aMC and the Madgraph on GPU development teams for our fruitful collaboration on these projects. In particular, I warmly thank Olivier Mattelaer for our initial discussions and our coding together for the new developments shown in this paper during the 2022 GPU Hackathon and more generally for our effective collaboration over many years on the Madgraph on GPU developments. This project would not have been possible without him. I am grateful to the organisers and our mentors at the CSCS GPU Hackathon in Lugano in September 2022, and I thank the other participants from our team, Stephan Hageböck, Stefan Roiser and Zenny Wettersten, for our useful discussions. I kindly acknowledge the use of LUMI HPC resources under project 465001592 (CERN / Madgraph GPU porting, EHPC-DEV-2024D11-007) to derive the results I showed for AMD GPUs.