### v2

# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

Abstract
The first production release of the CUDACPP plugin for the Madgraph5_aMC@NLO generator, which speeds up matrix element (ME) calculations for leading-order (LO) processes using a data parallel approach on vector CPUs and GPUs, was delivered in October 2024. This has been described in previous publications by the team behind that effort. In this paper, I describe my work on some additional developments and optimizations of CUDACPP, mainly but not exclusively for GPUs. The new approach, which represents a major restructuring of the CUDACPP computational engine, mainly consists in splitting the calculation of the ME, which was previously performed using a single large GPU kernel, into many smaller kernels. A first batch of changes, involving the move to separate “helicity streams” and the optional offloading of QCD color sums to BLAS, also via GPU tensor cores, was recently merged into a new production release of CUDACPP, in collaboration with my colleagues. Since then, I have completed a second batch of changes, which I now also consider ready to be merged in production, involving the possibility to split the calculation into groups of Feynman diagrams in separate source code files. This new feature makes it possible to compute QCD matrix elements for physics processes with a larger number of final state gluons: in particular, I present the first performance results from CUDACPP for the $2\!\rightarrow\!6$ process $gg\!\rightarrow\!t\bar{t}gggg$ on CPUs and GPUs and the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$ on CPUs, which involve over 15k and 230k Feynman diagrams, respectively. I also take this opportunity to describe in detail some features of the CUDACPP software which had not yet been documented, both in the GPU and vector CPU implementations.
journal: Eur. Phys. J. Cjournal: Eur. Phys. J. C
1 Introduction
The MadGraph5_aMC@NLO bib:mg5 ; bib:mg5amc physics event generator (in the following, MG5aMC) is an essential component of the data processing and analysis workflows of many high-energy physics (HEP) experiments, notably those at CERN’s Large Hadron Collider (LHC). Using Monte Carlo (MC) techniques, MG5aMC allows the calculation of cross sections and the generation of events to provide theoretical predictions against which experimental measurements can be compared. The MG5aMC software has been developed over more than two decades and does not yet fully exploit the potential of modern computing architectures. As is the case for other generators, this is a source of concern because event generation in the LHC experiments has a large computational cost, which is predicted to further increase during the High-Luminosity LHC (HL-LHC) programme bib:csbs .
In this context, a recent progress has been the port and optimization of the MG5aMC software for data parallel processing on graphical processing units (GPUs) and on CPUs with SIMD (Single Instruction Multiple Data) vector instructions. After a five-year development programme, which was initially facilitated by the activities of the HEP Software Foundation (HSF) event generator working group bib:csbs , and which regularly reported its progress through conference proceedings bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 , this project delivered its first production release of the MG5aMC code with GPU and SIMD support for QCD leading-order (LO) processes in October 2024. This has been reported at the CHEP2024 conference bib:chep2024 and in a more recent journal submission bib:zw2025 . MG5aMC is a code-generating framework, largely written in Python, which allows users to generate code, by default using Fortran, for any physics process of their choice. For complex processes, the computational bottleneck of the MG5aMC workflow is the calculation of the matrix element (ME) for each phase space point from the momenta of the initial and final state particles in that event. The main outcome of this project has been a new code-generating plugin, which was named “CUDACPP” because the generated physics code uses C++ instead of Fortran to execute the calculation of MEs on vector CPUs, and also includes CUDA extensions to run the code on NVidia GPUs (as well as HIP extensions for AMD GPUs). The new code speeds up the ME computation by performing it for many events at the same time using a data parallel approach bib:hsf2020 ; bib:lhcc2020 . To achieve this, work was also needed in the process-agnostic core MG5aMC event processing algorithm for LO processes, MadEvent, which is written in Fortran and with which the process-specific C++ code generated by CUDACPP is interfaced and linked: the main change is the move from a sequential single-event API to a parallel multi-event API, and the refactoring of some functions to make them stateless and re-entrant. This effort has been the result of a collaboration involving many existing and new contributors of the MG5aMC project, as detailed in the references cited above bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 .
In this paper, I describe my work on many additional enhancements and optimizations of CUDACPP, mainly but not exclusively targeting GPUs. These are developments for which the initial brainstorming with some of my colleagues from the development team, notably Olivier Mattelaer who prototyped with me the first concrete changes to the code, took place during the CSCS GPU Hackathon that we attended in Lugano in September 2022 bib:hack2022 . I then invested further effort in this area in February and especially in October 2024 bib:ks2024 after the first CUDACPP release, but had to pause this work again due to other constraints. In September 2025, I was finally able to resume this R&D activity, which I now consider essentially complete and which I am documenting in detail in this paper in view of its possible integration in new production versions of the CUDACPP software. A first batch of changes, which I had described in the preprint of a previous version of this paper bib:preprintV1 , has already been merged into a new production release of CUDACPP bib:cudacpp364 , in collaboration with my colleagues. Since then, I have completed a second batch of changes, which were still work in progress at the time of Ref. bib:preprintV1 , but which I now also consider ready to be merged in production.
1.1 Kernel splitting
The main idea behind all of these developments can be referred to as “kernel splitting”. In short, this consists in replacing the large monolithic kernel previously used by CUDACPP for the calculation of MEs from particle momenta, named sigmaKin, by many smaller kernels, executed either in parallel or sequentially, to achieve a more efficient and scalable use of the GPU. The aim is not only that of possibly achieving a higher peak throughput (in terms of MEs computed per second) with a large GPU grid, i.e. with a large number of events processed in parallel on the GPU during one offloading cycle, but also that of achieving a faster increase of throughput already with smaller GPU grids. This was the main focus of my first batch of changes, which have since been released. An important practical difference between the traditional MG5aMC workflow on CPUs and that using GPUs, in fact, is that the former was designed to handle iterations involving hundreds of events at a time, while the latter typically needs many thousands of events to be efficient: in this context, achieving higher throughputs with smaller grids would provide one way to use more manageable, but still reasonably efficient, event generation jobs on GPUs. An additional benefit of kernel splitting, which eventually became the main focus of my second batch of changes, is that smaller kernels are much more manageable by humans and compilers alike (both for CPUs and GPUs), and make it possible to achieve complex calculations that are instead very difficult if not impossible to perform with large monolithic kernels, such as the calculations of QCD matrix elements for physics processes with a large number of final state gluons. In this paper, in particular, I present the first performance results from CUDACPP for the $2\!\rightarrow\!6$ process $gg\!\rightarrow\!t\bar{t}gggg$ on CPUs and GPUs and for the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$ on CPUs, which involve over 15k and over 230k Feynman diagrams, respectively.
ihel0
cudacpp v1.00.02/3.6.3
one single kernel for all helicities (monolithic sigmaKin kernel)
ihel1
one kernel per helicity; one CUDA Stream per helicity
ihel2
two kernels per helicity: color sum & diagrams
QCD
color sum
Feynman
diagrams
ihel3
color sum: one kernel per helicity OR one BLAS per helicity
ihel4
one kernel per diagram in each helicity stream
ihel3p1
cudacpp v1.01.01/3.6.4
color sum: one kernel per helicity OR one single BLAS for all helicities
ihel4p1
one kernel per diagram in each helicity stream; CUDA Graph orchestration
ihel5
one kernel per diagram group in each helicity stream; separate source code files
ihel6p1
(DCDIAG=0) one kernel per diagram group in each helicity stream, exchange wavefunctions in GPU global memory, exchange color amplitudes in GPU global memory OR (DCDIAG=1) one single kernel for all diagram groups in each helicity stream, exchange wavefunctions in GPU registers, exchange color amplitudes in GPU registers
ihel6p2
remove templates from helicity amplitude methods
Figure 1: Schematic representation of the architectural evolution of matrix element calculations in the MG5aMC CUDACPP plugin, in the context of the kernel splitting developments presented in this paper. The steps above the blue dotted line, up to ihel3 and ihel4 included, have been presented in a previous preprint bib:preprintV1 of this paper, while those below it are described here for the first time. The steps in the green boxes, up to ihel3p1 included, have been merged upstream and included in production releases of CUDACPP; the steps in the red boxes, conversely, from ihel4 to ihel6p2 included, are included in a pull request that I recommend for merging upstream. More details about each step are provided in the text and in other figures.
Schematically, in this paper I describe my work on six sets of kernel splitting developments: (1) helicity streams; (2) color sum splitting into separate GPU kernels; (3) color sum host refactoring with optional BLAS offloading; (4) Feynman diagram splitting, with one separate GPU kernel for each diagram; (5) Feynman diagram splitting, with one GPU kernel for each group of diagrams, and separate source code files; (6) Feynman diagram splitting, with a single GPU kernel processing all diagram groups as GPU device functions. A brief description of these steps and of the connection between them is shown in Fig. 1. The performance of each set of developments was tested for different physics processes, software configurations and hardware architectures. While these developments were initially aimed at improving throughputs on GPUs, the large code refactoring that they imply was also propagated to the vectorised C++ code for SIMD CPUs, as code generated by the CUDACPP plugin is strictly the same for CPUs and GPUs, and the distinction between the two cases is only done at build time through #ifdef directives. I therefore analysed not only CUDA and HIP performance on NVidia and AMD GPUs, but also C++ performance on vector CPUs using different SIMD configurations. The possibility to split the ME calculation into groups of Feynman diagrams in different files, in particular, was the main enabling factor for being able to compute complex $2\!\rightarrow\!7$ physics processes on CPUs.
Considering the results of all my tests, my recommendation is to include the full set of these enhancements in new production releases of CUDACPP. This differs from the situation that I had described in my previous preprint bib:preprintV1 , where the diagram splitting work had not yet been completed. A first batch of changes (green boxes in Fig. 1) has already been merged into a new production release, in collaboration with my colleagues: this includes the first three steps of this work, notably helicity streams, the separation of color sums and Feynman diagrams, and the optional BLAS offloading of color sums (which was significantly improved with respect to that described in Ref. bib:preprintV1 ). The second batch of developments (red boxes in Fig. 1), notably including the splitting of Feynman diagrams into different groups and source code files, is in my opinion now also ready to be merged upstream. I note, in particular, that the performance of the code in its default configuration would not be negatively impacted by these changes: many of the features that I have added, in fact, are optional and are only enabled with specific flags at generation time, build time or run time (this includes the BLAS offloading of color sums, the ME splitting into groups of Feynman diagrams, their orchestration via CUDA graphs and the choice to treat diagram groups as separate kernels or as device functions within a single kernel).
1.2 Outline of this paper
As these new developments imply a significant refactoring of the existing CUDACPP codebase, and rely on internal features that have not yet been documented, I also found it useful to start this article with a brief review of some architectural aspects of the software, focusing on those most relevant to the kernel splitting developments, but also covering some details of the SIMD implementation. In this context, I should mention that the internal design of the CUDACPP plugin, which was largely my own work, is far from perfect and the code could certainly benefit from an extensive cleanup. In particular, over time I have added various switches to support alternative implementations of some software components, as I have done again in many of the new developments described in this paper: I believe that this has been very useful, and in some cases essential, to try out different approaches and eventually converge on well-defined solutions to address specific problems, but I am also aware that this results in code that might be more difficult to read, and where some of this flexibility may be regarded as an unnecessary complication. I hope that this article may help clarify the rationale behind some of these design choices and provide some guidance for possible future developments and/or cleanups. I stress, in any case, that the opinions that I express in this paper are only my own, and may differ from those of some of my colleagues.
The outline of this paper is the following. In Sec. 2, I briefly review some aspects of the CUDACPP software architecture design, focusing on those relevant to GPU kernel splitting, but also covering some details of the SIMD implementation for vector CPUs. In Sec. 3, I describe in detail the software design and implementation of my first batch of kernel splitting enhancements, and I provide the results and an analysis of various performance tests: this includes and partially extends the material that I had covered in my previous preprint bib:preprintV1 , with the exception of Feynman diagram splitting. In Sec. 4, I describe the various options I introduced to split the calculation of Feynman diagrams into different groups of diagrams and/or different kernels, and different source code files; I also give the first CUDACPP performance results for physics processes with large numbers of final state gluons, which could be computed thanks to these new features. In Sec. 5, I finally give my conclusions and an an outlook for this work.
Figure 2: Schematic representation of the architectural evolution of the work on the MG5aMC CUDACPP plugin between 2020 and 2024. This plot was prepared for CHEP2024 and is taken as-is from its proceedings bib:chep2024 , where additional details can be found.
2 Software architecture review
The overall software architecture design of the CUDACPP plugin for GPUs and vector CPUs, its evolution over time, its interplay with the MG5aMC framework and many specific implementation choices have already been described in the previous papers bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 from our team and I will not repeat this here. The main point of this section, instead, is to give further information about some software aspects which have not yet been documented, mainly those relevant to GPU kernel splitting, such as memory layouts and memory access, but also some related to the SIMD implementation for vector CPUs. First, in any case, I will give a very high level overview of the various areas and phases of development on CUDACPP, to put into context the new kernel splitting developments.
2.1 Three software areas and development phases
The evolution of the work on the CUDACPP plugin between 2020 and 2024 is represented schematically in Fig. 2 and was summarised in detail in the CHEP2024 proceedings bib:chep2024 , from which that plot is taken. Many more details about the progress of this work over time, including snapshots and links to the initial presentations of some of these ideas at working group meetings of the Madgraph on GPU project, can be found in the slides of my CHEP2024 presentation bib:chep2024slides .
The main aim of the CUDACPP software, from the very beginning, has been to use event-level data parallelism to try and speed up the calculation of MEs, which is the computational bottleneck of the whole MG5aMC framework. This is because, to a large extent, the same mathematical functions need to be numerically computed for different events during the ME calculation, and this makes it possible to efficiently compute the MEs for many events in lockstep at the same time using GPUs and SIMD CPUs. This is something that I had pointed out in two presentations to the HSF generator working group bib:hsf2020 and the LHCC bib:lhcc2020 in 2020, using plots from which I later derived those in Fig. 2.
Largely speaking, the work on the CUDACPP plugin has concerned three areas of the software, roughly corresponding to three phases of development:
  1. 1.
    ME engines: the port to GPUs and SIMD CPUs, using CUDA/HIP and vectorised C++, of the computational engines to calculate MEs from particle momenta, and their backport to the code-generating Python framework, for different physics processes.
  2. 2.
    MadEvent integration: the injection of the new code into one madevent application, by modifying the existing Fortran and linking the new ME engine as a C++ library into it.
  3. 3.
    Full workflow orchestration: the integration and testing of the full MG5aMC workflow involving many madevent applications, including the packaging and installation of the CUDACPP plugin.
Over time, the focus of developments has gradually shifted from the first to the last of these software areas, even if some work on ME engines and on the integration of C++ and Fortran continued up until the release in October 2024. This is reflected in the documentation provided in our previous papers. What I want to stress here, to put it into context, is that the new kernel splitting work that I describe in this paper was all done at the level of the ME engines alone, even if its motivation (as is the case for all of the work in the Madgraph on GPU project) lies in the streamlining of the full MG5aMC workflow. I mention this also because I believe that, amongst the many previous documents from this project, those describing the software features most relevant to kernel splitting in the ME engines are actually the very first ones, namely the proceedings bib:vchep2021 and presentation bib:vchep2021video from the vCHEP2021 conference.
More specifically, in all of the work that I present in this paper, I only developed code using the standalone application check.exe, where all three main components of an event generation application (random numbers, phase space sampling and ME calculation) are implemented in C++ and CUDA using a data parallel approach. This is because check.exe makes it possible to focus on the optimization of the ME engine without being constrained by the MadEvent Fortran infrastructure, where the legacy phase space sampling and other sequential non-ME components slow down the whole workflow. I routinely built check.exe using the CUDACPP makefile cudacpp.mk, without even building MadEvent or any other Fortran code (which proceeds through a separate makefile that internally delegates the build of the CUDACPP library to cudacpp.mk).
int main(…) // application check.exe (similar code in class Bridge)
int nevt = m_gpuBlocks * m_gpuThreads; // number of events == GPU grid size
DeviceBufferMomenta m_devMomenta( nevt ); // memory buffer for nevt events
DeviceBufferMatrixElements m_devMEs( nevt ); // memory buffer for nevt events
// Compute MEs for nevt events (m_gpuBlocks * m_gpuThreads) in one go
MatrixElementKernelDevice mek( m_devMomenta, m_devMEs, m_gpublocks, m_gputhreads, );
mek.computeMatrixElements( );
void MatrixElementKernelDevice::computeMatrixElements( )
gpuLaunchKernel( sigmaKin, m_gpuBlocks, m_gpuThreads, m_momenta.data(), m_MEs.data(), );
__global__ void
sigmaKin( const fptype* momenta, // input: momenta[nevt*npar*4]
fptype* MEs, ) // output: MEs[nevt], final sum over ihel
const int ievt = blockDim.x * blockIdx.x + threadIdx.x; // one event <-> one thread
// Zero the running sum MEs[ievt] and add the contribution of each helicity ihel
MEs[ievt] = 0;
for( int ihel = )
calculate_wavefunctions( ihel, momenta, MEs, );
// Randomly select a color and a helicity for event ievt
selcol[ievt] = …; selhel[ievt] = // extra output (from extra input random numbers)
__device__ void
calculate_wavefunctions( int ihel,
const fptype* momenta, // input: momenta[nevt*npar*4]
fptype* MEs, ) // output: MEs[nevt], running sum over ihel
using M_ACCESS = DeviceAccessMomenta; // non-trivial access (nevt events)
using E_ACCESS = DeviceAccessMatrixElements; // non-trivial access (nevt events)
using W_ACCESS = DeviceAccessWavefunctions; // trivial access (one event)
using A_ACCESS = DeviceAccessAmplitudes; // trivial access (one event)
fptype* wf[nwf] = …; // local variable for one event (wavefunctions)
cxtype amp[1]; // local variable for one event (amplitude for one Feynman diagram)
cxtype jamp[ncolor] = {}; // local variable for one event (dual amplitudes for color flows)
// Feynman diagram 1 of 2
vxxxxx<M_ACCESS, W_ACCESS>( momenta, ihel, …, wf[0] ); // compute wf[0]
vxxxxx<M_ACCESS, W_ACCESS>( momenta, ihel, …, wf[1] ); // compute wf[1]
oxxxxx<M_ACCESS, W_ACCESS>( momenta, ihel, …, wf[2] ); // compute wf[2]
ixxxxx<M_ACCESS, W_ACCESS>( momenta, ihel, …, wf[3] ); // compute wf[3]
VVV1P0_1<W_ACCESS, …>( wf[0], wf[1], …, wf[4] ); // compute wf[4]
FFV1_0<W_ACCESS, A_ACCESS, …>( wf[3], wf[2], wf[4], &amp[0] ); // compute amp[0]
jamp[…] += * amp[0];
// Feynman diagram N of N
// compute wavefunctions and amplitudes…
// Add ME contribution for helicity ihel (from quadratic form on color amplitudes)
fptype& ME = E_ACCESS::kernelAccess( MEs ); // ME for event ievt (from threadIdx.x etc)
ME += sum_ij ( cxconj((cxtype2)(jamp[i])) * colormatrix[i][j] * (cxtype2)(jamp[j]) );
Table 1: Pseudo-code of the ME calculation in CUDACPP before any kernel splitting changes (“ihel0”). Function main represents the check.exe standalone application. Note that sigmaKin is the only GPU kernel (__global__ function) in this workflow. Only a simplified GPU version of the code is shown, using scalar fptype floating point and cxtype complex number types: the actual code that supports both GPUs and SIMD CPUs uses fptype_sv and cxtype_sv data types inside calculate_wavefunction, making it possible to use the same formal code for scalar data and for SIMD vectors (see Ref. bib:vchep2021 for details). The CPU branch of the code for main, encapsulated by an #ifdef and not shown here, uses the Host versions of various classes instead of the Device versions. The pseudo-code for the color sum schematically indicates that this proceeds via a quadratic form using a color matrix, and that color amplitudes are converted from fptype/cxtype floating point precision to a potentially different fptype2/cxtype2 precision (this enables the “mixed” precision mode, where the former is based on double and the latter on float).
2.2 A closer look at the internals of the ME engine
Consider a physics process where npar is the total number of initial and final state particles. Assume that the calculation of Feynman diagram amplitudes is performed using floating point fptype, which is a typedef for either double or float. Very schematically, as shown in Fig. 2, the ME computational engine of CUDACPP is simply a software component that takes as input an array of particle momenta for many events (i.e. 4*npar fptype values for each of nevt events), and returns as its output an array of matrix elements (i.e. one fptype value for each of nevt events). The calculation in CUDACPP is actually more complex than this, as the ME engine also receives other inputs, such as: an event-by-event running coupling $\alpha_{s}$ (at a scale which is computed from particle momenta using an external Fortran module); two event-by-event random numbers for the selection of one specific helicity and leading color in the generated event (which are also returned as outputs); and an event-by-event channel identifier (for MadEvent single-diagram enhancement, also resulting in additional outputs). However, I will neglect this in the following, except otherwise stated, and focus only on the calculation of output MEs from input momenta.
In the production version of the CUDACPP plugin before any kernel splitting changes (hereafter “ihel0”), v1.00.02 in MG5aMC v3.6.3, the ME computational engine on GPUs is essentially a single and very large monolithic GPU kernel, named sigmaKin. This performs all of the relevant operations that are needed to compute the output ME from the input momenta for each event (one notable exception being the calculation of other couplings that depend on $\alpha_{s}$, which is delegated to a separate kernel). Very schematically, this is represented in the pseudo-code in Table 1. In particular, sigmaKin performs the following operations for each event: it keeps a running sum of the ME over helicities; it loops on all possible helicities of the external particles (or, more precisely, only on a pre-determined set of “good” helicities whose contribution is non-zero); it adds to the running sum of the ME the contribution from each helicity, by calling a device function calculate_wavefunctions which internally involves for each QCD color flow the calculation of particle and propagator wavefunctions and that of dual amplitudes, followed by the sum over all color flows using a quadratic form based on a color matrix; after the end of the helicity loop, sigmaKin randomly draws a helicity and a color for the generated event based on the ME contributions from individual helicities and colors. The calculation of helicity amplitudes for Feynman diagrams, in particular, is described in detail in Ref. bib:vchep2021 (see Fig. 1 therein); complementary details are also given in Ref. bib:zw2025 . As shown schematically in Table 1, this involves three types of elements to compute: the wavefunctions of external (initial/final state) particles, via functions like VXXXXX; the wavefunctions of internal propagators, e.g. via VVV1P0_1; the dual amplitudes for a given colour flow (“jamps”), e.g. via FFV1_0. The names and roles of these functions in CUDACPP are the same as in the original HELAS bib:helas and ALOHA bib:aloha implementations.
One key aspect in the design of this software chain, which has not been documented in detail so far, is the allocation of memory buffers and their access and use in computational kernels, in both the GPU and SIMD CPU code. This was a core part of my design and implementation work in the second half of 2020, which was motivated mainly by the aim of achieving SIMD speedups on CPUs, but also by the aim of achieving coalesced memory access on GPUs, and turned out to be a key enabling factor for both. The design principle which I adopted, in particular, is a complete separation of three key elements of the software (for details, see the backup slides 70–75 from December 2020 in Ref. bib:chep2024slides ):
  1. 1.
    Memory allocation. In the check.exe standalone application (or in the Bridge component that gets linked with madevent), CUDACPP allocates host and device memory buffers that are properly dimensioned for the number nevt of events that will be processed in parallel in one offloading cycle. On the GPU, nevt is simply the GPU grid size where kernels are launched (i.e. the product of the number of threads per block and of the number of blocks), as CUDACPP kernels use event-level data parallelism where each GPU thread processes one event. The classes responsible to allocate and hold the pointers to the memory buffers have no knowledge of the internal memory layout of the buffers; for instance, DeviceBufferMomenta in Table 1 allocates 4*npar*nevt fptype values, but it does not know the layout of the particle momenta arrays. In particular, as opposed to earlier versions of the code that were using structured memory allocations like std::vector in C++ and cudaMalloc3D in CUDA, from the end of 2020 the code uses unstructured memory allocations of raw buffers, both on the host (via malloc) for the GPU and SIMD CPU code, and on the device (via cudaMalloc) for the GPU code.
  2. 2.
    Memory layout and data access. The specific layout chosen for storing data inside the raw memory buffers allocated in the previous step is encapsulated in a separate set of memory access classes. All these classes have methods named kernelAccess, or some variation of this. On the GPU, these methods take as input the fptype* pointer associated to a raw memory buffer for nevt events, and they return as output a data item for the single event ievt processed by the GPU thread where the code is executed, typically indexed by the identifier of the GPU thread, blockDim.x*blockIdx.x+threadIdx.x. In some cases, these methods have additional parameters: method kernelAccessIp4IparConst in the DeviceAccessMomenta class, for instance, takes two additional parameters ip4 and ipar in order to retrieve a given 4-vector component for a given particle. It is only the memory access classes that, internally, are able to decode the memory layouts of a raw buffer: this is done by interpreting the buffers as one-dimensional C-style array or casting them as multi-dimensional C-style arrays. For most data buffers, a Structure-Of-Array (SOA) or an Array-Of-Structure-Of-Array (AOSOA) layout is chosen, where the data items of a given type for different events are contiguous: as explained in detail in the vCHEP2021 proceedings bib:vchep2021 , this is absolutely essential for SIMD processing on vector CPUs, and it is also beneficial – but not at all a strict requirement – on GPUs to improve performance through coalesced data access. In some cases, such as couplings, the floating point raw arrays include the real and imaginary components of complex data, and it is the responsibility of the data access classes to return outputs with the API of a complex number data type. The case of CPU code includes the additional complication of returning outputs whose data types fptype_sv and cxtype_sv can be scalar (for no-SIMD C++) or SIMD vectors fptype_v and cxtype_v (through compiler vector extensions bib:vchep2021 , hereafter CVEs) of fptype or cxtype data. Concerning the use of memory access classes in other CUDACPP software components, most notably in the computational methods for helicity amplitudes like VXXXXX, VVV1P0_1 or FFV1_0, I initially chose to implement this as a template parameter: in Table 1, for instance, DeviceAccessMomenta is a template parameter M_ACCESS of VXXXXX. The idea behind the choice of using templates was to allow the flexibility of easily switching between different memory layouts at build time, to compare their performances; in retrospective, the same flexibility may have been achieved with the same memory access classes but without templates [^1] As noted in Sec. 4, eventually I did decide to remove those templates, even if it was not clear to me a priori if this would make the code easier to read and maintain, or faster to build: my main motivation was an attempt to allow some CUDA builds in some configurations of particularly complex $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes, which were otherwise crashing or taking too long. Unfortunately, my tests about the effectiveness of template removal for simplifying CUDA builds were at best inconclusive, as those builds continued to fail; however, removing templates did not have any negative impact on performance elsewhere, and I suggest that this may represent a useful cleanup to merge upstream and integrate in a future production release. .
  3. 3.
    Arithmetic operations and other computational functions. Finally, the decoupling of data access from actual calculations has probably been one of the most important aspects of the whole CUDACPP software design. Just like the memory allocation classes, also the helicity amplitude functions like VXXXXX, VVV1P0_1 or FFV1_0 have almost [^2] One notable exception is that some assumptions are still made about the memory layout of wavefunctions, which for spin 1/2 and spin 1 particles are complex 6-dimensional arrays as in the original HELAS bib:helas and ALOHA bib:aloha implementations, from which CUDACPP is derived. This detail is relevant to the kernel splitting for different Feynman diagrams, as discussed in Sec. 4. no knowledge of the memory layouts used in the multi-event data buffers. These functions also ignore whether the calculations are performed in single or double precision, as I encapsulated this choice in the fptype type definition. Even more, these functions ignore whether the arithmetic operations within them are applied to scalar values on GPUs or CPUs or to SIMD vectors on CPUs: in the latter case, this is possible because simple operators like “+” are automatically understood by the compiler as vector operations when applied to CVE vectors of fptype, and are also implemented using SIMD CVE operations in their definition for vectors of custom types, notably the vectors of cxtype complex types. This design is extremely powerful because it has made it possible to use formally the same exact lines of software in the code-generated helicity amplitude functions like VVV1P0_1 or FFV1_0. I note in passing, however, that the functions computing wavefunctions for initial and final state particles, like IXXXXX, OXXXXX and VXXXXX, or their variations for massless or beam-collinear particles, are not code-generated from a model Lagrangian, but are instead hardcoded and often required particular care and successive iterations. Two complications that I had to address for SIMD code, in particular, are the following [^3] In this context, I find it important to stress that, in my experience, achieving a robust and performant CUDACPP implementation for SIMD CPUs was far from a trivial task, and was, in fact, much more complex than achieving it for GPUs. The difficulties for implementing IXXXXX-like functions are just two examples. I also already mentioned the fact that there are very strict constraints in the memory layouts for CPU SIMD, unlike those for GPUs; as described in Ref. bib:vchep2021 , in particular, this also implied the development of wrappers and adapters for vectors cxtype_v of complex numbers, as their real and imaginary parts must be stored as an SOA using two contiguous fptype_v arrays (RRRRIIII), rather than as an AOS (RIRIRIRI). Another example is the implementation of mixed floating point mode, where I added a complex mechanism to merge/split two SIMD vectors of double into/from one SIMD vector of float at the boundary between dual amplitude and color sum calculations, in order to use the widest possible SIMD vectors in both cases and maximize efficiency. More generally, in my opinion, it is fair to say that achieving SIMD speedups through vectorization is always much more complicated than achieving speedups on GPUs. Code acceleration on GPUs, in fact, can be achieved even without lockstep processing (at the cost of some inefficiency from warp divergence), or without optimized memory layouts (at the cost of some inefficiency from the lack of coalesced data access), simply because a GPU has thousands of threads. One good feature of CUDACPP is that, by focusing on achieving C++ vectorization on SIMD CPUs, this has also improved the efficiency of the code on GPUs. : first, unlike VVV1P0_1 or FFV1_0, the IXXXXX-like functions include some if/else branching, which I reimplemented using vector masks in the SIMD case in order to fully exploit data parallel speedups (even if, from a performance point of view, this is only important for simple processes with few Feynman diagrams); second, as I performed all my functional testing with Floating Point Exception traps (FPEs) enabled in order to develop more robust code, I came across some crashes caused by the interplay of compiler optimizations, SIMD CVEs and FPE traps, which I addressed using volatile keywords and which pushed me to develop a large set of functional tests specifically for these functions.
On top of the separation of the three software concerns above (data allocation, data access and arithmetic calculations), another important part of the design of the CUDACPP software was the implementation of kernel launching on GPUs and event loops on CPUs. My aim here was to keep as much as possible of the software logic and of the actual code identical for GPUs and CPUs, in order to simplify the iterative addition of new features and of bug fixes for both cases. In short, I addressed this by keeping also for SIMD CPUs the idea of a “grid” of events that are processed in one given iteration, and by explicitly subdividing this grid into SIMD event vectors and adding an explicit loop over them. Much more practically, and more importantly for the kernel splitting work described in the next sections, the link between memory allocations and the execution of the ME computational engines is provided by two different incarnations of a MatrixElementKernel (MEK) class, one for GPUs (Device) and one for CPUs (Host). A singleton instance of the appropriate MEK class is constructed and used inside check.exe for standalone tests, or inside the Bridge component that is linked with Fortran in madevent for the full MG5aMC workflow. As shown in Table 1, it is the MatrixElementKernelDevice class that internally launches the monolithic sigmaKin kernel in the current CUDACPP. This is the starting point for the enhancements presented in the following sections.
2.3 A brief note on SIMD build modes
As mentioned previously, the new developments that I present in this paper are mainly optimizations that target GPUs specifically, but they are also relevant and useful for the CUDACPP implementation on vector CPUs. In this paper, in particular, I also present results from the five different SIMD build modes of CUDACPP, namely none, sse4, avx2, 512y and 512z. While these have already been described to some extent in earlier publications from our team bib:vchep2021 ; bib:zw2025 , I find it useful to take this paper as an opportunity to give a few additional and previously undocumented details.
The SIMD modes of CUDACPP were initially introduced targeting Intel CPUs. My design strategy was to define separate build modes, which on every specific system would allow not only the fastest calculation compatible with the hardware specs, but also an evaluation of the speedup achieved through the SIMD implementation of CUDACPP in vectorised C++. Largely speaking: none does not use any SIMD extensions; sse4 uses SSE4.2 instructions and 128-bit xmm registers; avx2 uses AVX2 instructions and 256-bit ymm registers; 512y uses AVX512 instructions and 256-bit ymm registers; 512z uses AVX512 instructions and 512-bit zmm registers. The 512y mode may appear surprising: the rationale behind it is, quite simply, that it is around 10% faster than avx2 on CPUs with the AVX512 instruction set, but it is also faster than 512z on AVX512 CPUs with a single FMA unit (such as Intel Xeon Silver 4216 CPUs); conversely, it is slower than 512z on AVX512 CPUs with two FMA unit (such as Intel Xeon Gold 6326 CPUs). This was first reported for the CUDACPP software in Ref. bib:ichep2022 and has been confirmed in many tests since then.
More in detail, each build mode corresponds to a well-defined choice of two parameters, the CPU instruction set, and the length of CUDACPP SIMD vectors:
  1. 1.
    The CPU instruction set is configured via the -march flag of the compiler, in the cudacpp.mk makefile. First and foremost, this must be well-defined because an application compiled for a given instruction set would crash if it is executed on a hardware that does not support it (for instance, if it is compiled for AVX512 but executed on a host that does not support it). While many different levels and sublevels of vectorization have been defined over time, such as AVX, AVX2, SSE4.1 or SSE4.2, the choice I made was to only support a very limited subset of these in CUDACPP. In particular: none uses -march=x86-64, which disables all SIMD extensions; sse4 uses -march=nehalem, which supports SSE4.2 extensions with 128-bit xmm registers; avx2 uses -march=haswell, which supports AVX2 extensions with 256-bit ymm registers; 512y and 512z both use -march=skylake-avx512, which supports a specific set of AVX512 extensions with 512-bit zmm registers. In this context, I chose not to foresee a build mode based on -march=native, which would offer the greatest set of hardware features available on a CPU, but is somewhat poorly defined. This is similar to what is done in some LHC experiments: LHCb, for instance, uses well-defined software architecture builds based on x86_64 Micro-Architecture Feature Levels bib:phoronix such as x86_64-v2. In retrospective, indeed a better choice would have been to configure -march based on x86_64 Feature Levels, but these were not yet widely used when I developed the SIMD implementation of CUDACPP in 2020.
  2. 2.
    The SIMD vector length is the core of the CUDACPP SIMD implementation, where it is encapsulated by a preprocessor variable MGONGPU_CPPSIMD and an equivalent C++ variable neppV. As described more in detail in Ref. bib:vchep2021 , in fact, the basic strategy of the CUDACPP SIMD implementation is the definition, using compiler vector extensions, of vectors fptype_v of neppV floating point variables of type fptype. The five build modes are defined as follows: none uses neppV=1, sse4 uses neppV=2 and neppV=4 in double and single precision, respectively; avx2 and 512y both use neppV=4 and neppV=8 in double and single precision; 512z uses neppV=8 and neppV=16 in double and single precision.
The SIMD instruction set and vector length in the five CUDACPP build modes are, of course, tightly related. The choice of the appropriate MGONGPU_CPPSIMD and neppV depends on the compiler options defined in the cudacpp.mk makefile. In most cases, notably none, sse4 and avx2, this depends on the preprocessor variables defined by the specific -march choice. In the case of 512y and 512z, however, which use the same -march, an additional compiler flag -DMGONGPU_PVW512 informs CUDACPP that it should “prefer vector width 512” in the 512z mode, i.e. zmm registers. In addition, the -mprefer-vector-width=256 flag is passed in the 512y mode, to enforce the use of ymm registers alone, although this is probably unnecessary.
Finally, I note that the names of the CUDACPP SIMD build modes are presently used also on non-Intel CPUs. For instance, on ARM CPUs the sse4 mode uses NEON vector instructions and 128-bit registers; this is defined via preprocessor directives that are currently being reviewed. This simplifies the user interfaces of the software, but is clearly a language abuse: eventually, it may be appropriate to switch to a better terminology. Apart from a rationalization of build modes, in any case, the port to new architectures should not pose any big challenge. Some opportunities for larger SIMD speedups may arise if new hardware architectures with larger vector registers became available, although this seems unlikely at the moment. From the CUDACPP point of view, this would only involve the definition of new build modes [^4] As a proof of concept, I already prepared an implementation with the values of neppV that would be appropriate to 1024-bit and 2048-bit registers, and tested its functionality. This is currently of no practical utility, also because, not surprisingly, it is slower than the existing CUDACPP SIMD modes. with larger values of neppV, but the overall implementation based on compiler vector extensions would remain essentially the same.
3 Kernel splitting I: helicity streams, color sum
As discussed extensively above, a key parameter of the CUDACPP ME computational engine is the number nevt of events that are processed in parallel in one offloading cycle (in practice: in one call of the MEK computeMatrixElements function in Table 1). On a GPU, this is simply equal to the GPU grid size, i.e. to the product of the number of blocks gpuBlocks and of the number of threads per block gpuThreads. All relevant multi-event arrays, both in the Fortran madevent application and in the CUDACPP Bridge component (or in the check.exe standalone application), must be large enough to contain nevt events: the larger the GPU grid size, in particular, the larger the RAM footprint of the application (which can become very large bib:ichep2022 for madevent). This parameter can be configured at build time (and partly at runtime) in the madevent application, where it is referred to as VECSIZE bib:zw2025 , and is highly configurable in check.exe, where gpuBlocks and gpuThreads are defined independently at runtime.
Physics process $g\!\rightarrow\!t\bar{t} gg$$\rightarrow\!t\bar{t}g gg$$\rightarrow\!t\bar{t}gg gg$$\rightarrow\!t\bar{t}ggg gg$$\rightarrow\!t\bar{t}gggg gg$$g\!\rightarrow\!t\bar{t}ggggg$
Feynman diagrams 3 16 123 1240 15495 231280
SDE channels 3 15 105 945 N/A N/A
Leading colors 2 6 24 120 720 5040
Color matrix 2 x 2 6 x 6 24 x 24 120 x 120 720 x 720 5040 x 5040
Helicities 16 32 64 128 256 512
Wave functions 5 12 26 121 750 5834
Code generation codebase ihel6p2 ihel6p2 ihel6p2 ihel6p2 ihel6p2 ihel6p1
Code generation mode mad mad mad mad sa sa
Code generation d.p.g. any any any any $\leq$100 $\leq$1000
Code generation d.p.f. any any any any $\geq$100 $\geq$1000
Code generation time 2s 2s 4s 32s 13m 30h
Feynman diagram groups 1 1 1 1 155 232
Feynman diagram files 1 1 1 1 155 232
LOC (color_sum.cc) 438 442 460 556 1156 5.4k
LOC (diagramsXYZ.cc) 135 307 1986 27k 652k 18.7M
Size (color_sum.cc) 24 kB 24 kB 28 kB 92 kB 2.6 MB 137 MB
Size (diagramsXYZ.cc) 8 kB 16 kB 84 kB 1.0 MB 28.2 MB 721 MB
Build time (five C++ modes) 6s 6s 6s 23s 3m 55m
Build time (CUDA DCDIAG=0) 11s 11s 14s 6m 9m failed
Build time (CUDA DCDIAG=1) 13s 13s 15s 3m 6m failed
Table 2: Comparison of the computational complexity of the six physics processes considered in the tests described in this paper. The rows represent the following: (1) the physics process; (2) the number of Feynman diagrams; (3) the number of distinct SDE channels, i.e. of the diagrams used in the MadEvent single diagram enhancement algorithm; (4) the number of leading QCD color flows; (5) the size of the color matrix; (6) the number of combinations of helicities for initial and final state particles (for these six processes, this coincides with the number of good helicities with non-zero contributions to MEs); (7) the number of wavefunctions to be computed for all external (initial and final state) and internal (propagator) particles; (8) the codebase used for code generation; (9) the code generation mode, “mad” for full generation including Fortran madevent and “sa” for standalone only; (9-10) the maximum number of diagrams per group (dpg) and the minimum number of diagrams per file (dpf) during code generation; (12) the code generation time on a typical CPU; (13-14) the resulting numbers of Feynman diagram groups and source code files; (15-17) the total number of lines of code and total size in the CUDAPP generated color_sum.cc and the sum of all diagramsXYZ.cc files; (19-21) the time for a parallel (-j) build of the software with a cold cache (CCACHE_RECACHE=1), either for all five C++ modes on an Intel Xeon Gold 6326, or for CUDA in each of two DCDIAG modes on an AMD EPYC 7313. All of these numbers are needed to describe the complexity of the ME calculation in a single standalone (or madevent) application, which is the main focus of this paper; the number of SDE channels, in addition, is one of the factors that determines how many instances of a madevent application must be launched in a full event generation workflow. The number of helicities determines how many GPU streams are used, as of the ihel1 software. The number of leading colors determines the number of dual amplitudes that are computed using Feynman diagrams and used as input to the color sum; as of ihel2, these are stored into and retrieved from GPU global memory. The size of the color matrix determines the complexity of the color sum calculation (addressed with a separate GPU kernel as of ihel2, and optionally via BLAS as of ihel3). The number of wavefunctions is particularly relevant to the ihel4-ihel6 software versions, which handle their storage and retrieval in GPU registers and/or GPU global memory in subtly different ways.
As I briefly hinted above, one of the main aims of the new developments presented in this section is the fact that the ME calculation throughput (in MEs/s) on a GPU is generally quite low for small grids i.e. small values of nevt, and only reaches a peak plateau for relatively large grids, i.e. large values of nevt. This is an issue that exists in CUDACPP since the very beginning of our developments: in Fig. 5 of our vCHEP2021 proceedings bib:vchep2021 ), for instance, we had shown that the ME throughput for the $gg\!\rightarrow\!t\bar{t}gg$ process only ramps up significantly with at least 16k events in the grid, and reaches the peak plateau even later, with 128k events in the grid. In the same paper, we had also already suggested that higher throughputs, possibly because of lower “register pressure”, might be achieved by splitting monolithic sigmaKin kernel into smaller kernels: two specific ideas that we had mentioned, in particular, were the possible use of different GPU threads to process different helicities in the same event, or the possible use of CUDA graphs to orchestrate a much larger number of smaller GPU kernels. I will come back to both of these ideas in the following.
Since the ramp-up of ME throughputs with increasing GPU grid sizes is one of the main issues that this new work aims at addressing, or in any case a very good test of the effectiveness of these code changes, the results that I present in this paper will mainly consist of plots of that sort. One limitation of this work is that, due to lack of time, I will only show plots for a single CPU process, and only for the standalone application check.exe. This is unfortunate, because the practical benefits of any progress in this area would mainly be for the full MG5aMC workflow, using several madevent applications accessing the GPU in parallel. In the ACAT2022 presentation bib:acat2022 (see Fig. 1 therein), I had also studied the ramp-up of throughput as a function of the GPU grid size when several check.exe applications are launched in parallel, which seemed to indicate that the use of the GPU is more efficient in that case. Those results had been obtained using software containers prepared for the HEP-SCORE benchmarking project bib:bmk ; it would be useful to repeat similar tests in the future, when updated HEP-SCORE containers are prepared using more recent versions of CUDACPP, possibly including those presented in this paper.
Concerning my software development process, I note that for this new research I used the same methodology that I have been following for all of my work on MG5aMC during the last four years, since the completion bib:ichep2022 of a fully functional CUDAPP code-generating plugin in the “epochX” development workflow at the end of 2021. In particular, I always prototyped, tested, fixed and optimized my changes using as baseline the code-generated software for one specific physics process, typically $gg\!\rightarrow\!t\bar{t}$ or $gg\!\rightarrow\!t\bar{t}gg$. In the case of this work on kernel splitting, I focused on the CUDA implementation but also tested the SIMD C++ version. I then committed all changes for the given process to git, until these reached a state that I considered reasonably complete. At that point, I backported my changes to the code-generating Python framework, and regenerated all physics processes. I then performed larger-scale functional and performance tests, also using different hardware implementations like AMD GPUs, and iterated until completion. This development process has been possible only because, while the git repository of MG5aMC bib:mg5amc-github mainly contains the hardcoded components of the framework and its code-generating engine, the git repository of the madgraph4gpu project bib:mg4gpu-github contains not only the CUDACPP plugin but also the code-generated software for several physics processes. In my experience, this has been a key ingredient in the development of these latest enhancements to CUDACPP, but also much more generally of the whole plugin. In particular, I stress that I would never have been able to design and implement such large code changes directly in the code-generating Python code.
Figure 3: Schematic representation of the CUDACPP engine for computing MEs (sigmaKin), and of its evolution through the first four scenarios described in this paper: (ihel0) current version before kernel splitting; (ihel1) helicity streams; (ihel2) color sum kernel; (ihel3b) color sum on BLAS via host dispatcher. For the ihel3 software, only the (non-default) case with BLAS enabled at runtime is illustrated: by default, the ihel3 software has BLAS disabled at runtime, which is essentially the same as what is shown for the ihel2 scenario (the only difference is that in the ihel3 scenario the kernel is named color_sum_kernel and is invoked by a color_sum_gpu host function, which could also dispatch the calculation to the color_sum_blas BLAS host function).
In the following subsections, I will describe the first batch of three sets of kernel splitting changes that I developed, and which have already been integrated in a new production release of CUDACPP: (ihel1) helicity streams; (ihel2) color sum splitting into separate GPU kernels; (ihel3) color sum host refactoring with optional BLAS offloading. As shown in Fig. 1, these developments are sequential, i.e. ihel3 includes ihel2, which includes ihel1. The software architecture of the sigmaKin ME engine for these three scenarios is represented in Fig. 3, where it is also compared to the situation before any kernel splitting, ihel0. Note that I will not provide new pseudo-code listings for these kernel splitting developments, but in some cases I will refer to the pseudo-code for version ihel0 in Table 1 to point out what I changed or which technical issues I had to address. The results of my tests for these different versions of the software, as well as for those in my second batch of changes described in the next section, are given in Fig. 4 for an NVidia V100 GPU at CERN, in Fig. 5 for an AMD MI200 GPU at LUMI, and in Fig. 6 for an Intel Xeon Silver 4216 CPU, for three floating point precisions (double, mixed, float) and for four physics processes of increasing complexity ($g\!\rightarrow\!t\bar{t}, gg$$\rightarrow\!t\bar{t}g, gg$$\rightarrow\!t\bar{t}gg, gg$$g\!\rightarrow\!t\bar{t}ggg$), whose relevant parameters are described in detail in Table 2.
Figure 4: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU (node itscrd90 at CERN). Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. The number of threads per block is fixed to 32 (NVidia GPU warp size); the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel1) helicity streams; (ihel3) color sum kernel; (ihel3b) cuBLAS color sum; (ihel3p1) release v1.01.01, color sum kernel; (ihel3b) release v1.01.01, cuBLAS color sum; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, color sum kernel; (ihel6p2b) diagram groups, cuBLAS color sum. For ihel6p2 and ihel6p2b, all four processes were generated with a single diagram group, which was executed as a kernel at runtime, without graphs.
Figure 5: Throughputs (ME/s) as a function of grid size for an AMD MI200 GPU (at LUMI). Higher is better. The 9 plots correspond to 3 physics processes in 3 floating point precisions. The number of threads per block is fixed to 256; the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel1) helicity streams; (ihel3) color sum kernel; (ihel3b) hipBLAS color sum; (ihel3p1) release v1.01.01, color sum kernel; (ihel3b) release v1.01.01, hipBLAS color sum; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, color sum kernel; (ihel6p2b) diagram groups, hipBLAS color sum. For ihel6p2 and ihel6p2b, all four processes were generated with a single diagram group, which was executed as a kernel at runtime, without graphs.
Figure 6: Throughputs (ME/s) as a function of CUDACPP SIMD build mode for an Intel Xeon Silver 4216 CPU (node itscrd90 at CERN). Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel3p1) release v1.01.01; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, using code generated with a single diagram group.
3.1 Helicity streams (“ihel1”)
The internal substructure of the sigmaKin ME engine in the previous version of CUDACPP before my kernel splitting changes (ihel0) is illustrated schematically in the top-left diagram in Fig. 3. This gives a visual representation of the pseudo-code in Table 1. As mentioned previously and as visible in the diagram, the two main components of the calculation for each event, namely the computation of wavefunctions and dual amplitudes for a given color flow from Feynman diagrams, and their squaring and sum over all color flows, are performed sequentially for all helicities. This workflow makes it impossible to split these two components into separate kernels, because the loop over helicities is effectively inside the loop over events.
The very first step in my kernel splitting developments was therefore, quite naturally, to reverse this situation and make the event loop the innermost loop, inside an outermost loop on helicities. This was the focus of my work with Olivier Mattelaer during the 2022 GPU hackathon (and is quite likely an idea that he suggested, in fact). In practice, the main change to achieve here was to turn sigmaKin from a __global__ device kernel into a host function, and to turn instead calculate_wavefunction from a __device__ function callable by a kernel into a kernel itself. Other computations also had to be modified: for instance, the selections of event-by-event colors and helicities, which the sigmaKin kernel was performing outside the helicity loop (see Table 1) have now also been turned into separate GPU kernels. As discussed in Ref. bib:ks2024 , by themselves these changes already provide a moderate increase of throughputs for small grids.
The real breakthrough, however, came when the parallel calculations for different helicities were moved to separate CUDA Streams. In this “ihel1” version of the software, which is illustrated in the top-right diagram in Fig. 3, ME throughputs on NVidia GPUs reach their peak performance with much smaller grids than in the previous ihel0 version, for all physics processes and floating precisions I tested. The peak throughput themselves are also increased by around 10-20%. This can be seen in Fig. 4, by comparing the blue (ihel0) and orange (ihel1) curves. The improvement is especially impressive for complex processes like $gg\!\rightarrow\!t\bar{t}ggg$, where peak throughputs are reached with $\cal O$}(100) events per grid in ihel1, as opposed to $\cal O$}(10k) in the current ihel0. These improvements from ihel1 are essentially the single most important progress described in this paper. Most likely, the improvement comes from the increase in parallelism of the workflow: instead of a single kernel launch in each offloading cycle, which runs for a long time because it internally loops over helicities, there are now several, much shorter kernels launched in parallel, one for each helicity, in a separate GPU stream for each helicity.
For AMD GPUs, where the same solution was implemented using HIP Streams, the benefits are much less clear, as shown in Fig. 5: throughputs increase with both small and large grids for complex processes like $gg\!\rightarrow\!t\bar{t}gg$, but for simpler processes the opposite effect is observed [^5] I show no results for $gg\!\rightarrow\!t\bar{t}ggg$ on AMD GPUs, because the code failed to build on the system I used (“error: unhandled SGPR spill to memory”) with the ihel0 software. It is possible that this issue could be addressed using the Feynman diagram splitting techniques in ihel5-ihel6, but I did not attempt this. More generally, I stress that, while in this paper I present results also for AMD GPUs, I never attempted any serious optimization targeting them specifically, and it is perfectly possible that CUDACPP performance on this platform could be significantly improved with relatively little effort. . The code refactoring in ihel1 was also propagated to the C++ code, and tested on a reference Intel CPU; these showed that the throughputs are essentially unchanged for all SIMD builds.
3.2 Color sum as a separate GPU kernel (“ihel2”)
The next logical step in my kernel splitting developments consisted in separating the two main components of the ME calculation, namely, (1) the calculation from Feynman diagrams of the dual amplitudes for ncolor color flows (jamp: a vector $J$ with ncolor complex elements) and (2) the color sum. Initially, I simply split the calculate_wavefunction kernel, which was doing both computations in ihel1, into two separate kernels: calculate_jamps, which calculates the jamp $J$, and color_sum, which computes the quadratic form $J^{H}(C)J$ for a symmetric real color matrix $C$ using the vector $J$ and its conjugate transpose $J^{H}$. This “ihel2” version of the software is illustrated in the bottom-left diagram in Fig. 3. One important difference in this case is that the jamp variable is no longer a local variable inside calculate_wavefunction, as it was in ihel0 (see Table 1) and ihel1: instead, it is now a GPU global memory buffer that is allocated outside the MEK component and is accessed, with the appropriate layout decoding provided by a new class DeviceAccessJamps, in both the calculate_jamps and color_sum kernels.
As the plots I prepared are already packed, I decided not to present the ihel2 results explicitly: for reasons that will become clear in the following, in fact, these are essentially identical to the “ihel3” results shown by the solid-red curve in Fig. 4. As can be seen by comparing this to the orange curve for ihel1, on Nvidia GPUs the further change in the software from the split of Feynman diagrams and color sums yields an additional increase of the peak throughput by 10-20% for complex processes like $gg\!\rightarrow\!t\bar{t}gg$ and $gg\!\rightarrow\!t\bar{t}ggg$. For simpler processes, conversely, it results in a minor decrease of peak throughputs; in my opinion, this is a moderate cost that can be tolerated, as speeding up complex processes is more important. On AMD GPUs, as shown in Fig. 5, the additional change results in a minor increase in peak throughputs for $gg\!\rightarrow\!t\bar{t}gg$, but for simpler processes ihel2 is almost indistinguishable from ihel1. In the CPU implementation, throughputs are again essentially unchanged for all SIMD builds.
3.3 Color sum dispatcher to kernel or BLAS (“ihel3”)
The next step of my developments consisted in investigating the possible use of the cuBLAS linear algebra library for computing color sums on GPUs, instead of using a CUDA kernel. One of the main motivations for this work was that the current CUDACPP code only uses the traditional CUDA cores, but a large part of the computing power on recent NVidia GPUs comes from specialized tensor cores designed for the matrix algebra operations used in AI, and cuBLAS may provide a way for CUDACPP to exploit them. Developing code for tensor cores, in fact, is challenging because it requires the use of programming APIs other than CUDA: an easier alternative consists in using specialized libraries for AI or linear algebra that internally use the tensor core APIs, cuBLAS being one of them.
The BLAS implementation I developed is fully integrated in CUDACPP, and its functionality has been extensively tested both in the standalone use case and in the full MG5aMC workflows. It was originally developed for cuBLAS on Nvidia GPUs, but it has also been ported to AMD GPUs using the hipBLAS wrapper for rocBLAS. Studies of standalone color sums with cuBLAS and tensor cores had already been done bib:srblas in previous years by my colleagues in the Madgraph on GPU project: the work that I present in this section, however, is not based on the code developed for those studies and represents a restart from first principles.
Figure 7: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine, for different physics processes and software configurations, in the ihel3 version of the software. Lower is better. The 3 plots correspond to 3 floating point precisions. Measurements performed on a node equipped with an NVidia V100 GPU and an Intel Xeon Silver 4216 CPU (which supports AVX512 but gives better results for CUDACPP in the “512y” mode than in the “512z” mode because it has only one FMA unit). The results shown for CUDA are based on the default ihel3 implementation of color sums using GPU kernels.
In practice, my work on this “ihel3” version of the software was the following. To start with, I encapsulated the color sum calculation on GPUs in a host function color_sum_gpu (in parallel, I also created a color_sum_cpu function for the vectorized C++ version on SIMD CPUs). To make the software more modular and more manageable, I also took this opportunity to clean up the color sum code and move it to a separate source code file. The color_sum_gpu host function is just a wrapper that may dispatch the color sum calculation to two different GPU implementations:
  1. 1.
    color_sum_kernel. By default, the calculation is performed using a kernel color_sum_kernel. This is essentially the same code as in the ihel2 software, with one minor difference: the GPU global memory layout of the dual amplitudes, which was an SOA jamp[ncolor][2][nevt] in ihel2, is now an SOA jamp[2][ncolor][nevt], because this makes it easier to separate the real and imaginary parts of the dual amplitudes for the BLAS calculation, and the same layout is used for simplicity in both the kernel and BLAS implementations. The achieved performance is represented by the “ihel3” solid-red curve in Fig. 4. As mentioned above, these results are indistinguishable from those of ihel2, which I therefore decided not to show this explicitly to simplify the plot; the comparison, however, is available in my previous preprint of this article bib:preprintV1 .
  2. 2.
    color_sum_blas. The second implementation of the color sum consists in a host function color_sum_blas that internally calls the BLAS library. Specifically, since the color matrix $C$ is real and symmetric, the color sum $J^{H}(C)J$ over the vector of dual amplitudes $J=A+iB$ may be decomposed as
    $(A^{t}-iB^{t})(C)(A+iB)=A^{t}(C)A+B^{t}(C)B,$ (1)
    i.e. as the sum of two quadratic forms $V^{t}(C)V$, where the real vector $V$ may represent either the real part $A$ or the imaginary part $B$ of the complex dual amplitude vector $J$. Each calculation involves two steps. In mixed or float precision, for instance, the vector $(C)V$ is computed using cublasSgemm, while its dot product with $V^{t}$ is then computed using cublasSgemmStridedBatched. These two functions, as well as their double-precision and their HIP counterparts, are called through their abstractions via #define directives, in the header-only approach described in Ref. bib:chep2024 . The intermediate results of the first calculation, $(C)V$, are stored in GPU global memory using an additional buffer allocated in the MEK component (the allocation is done at runtime after determining the number of “good” helicities in each physics process, since the amount of memory allocated is proportional to the number of good helicities). As shown in the bottom-right diagram in Fig. 3, the BLAS calculations for different helicities are performed in separate GPU streams: technically, many BLAS handles are used, each associated to a different stream. Since the performance of the BLAS implementation of the color sum is generally worse than that using kernels, as discussed below, this is only available as an option, which must be explicitly enabled at runtime by setting an environment variable [^6] Set environment variable CUDACPP_RUNTIME_BLASCOLORSUM at runtime to use the cuBLAS (on Nvidia GPUs) or hipBLAS (on AMD GPUs) implementation of color sums.. Finally, as the BLAS library contains several switches targeting tensor cores, I also added another environment variable [^7] Set environment variable CUDACPP_RUNTIME_CUBLASTF32TENSOR at runtime to encourage cuBLAS to use tensor cores in color sums. This sets math mode CUBLAS_TF32_TENSOR_OP_MATH. to encourage BLAS to use tensor cores in the color sum (TF32 math mode).
The performances of the cuBLAS and kernel implementation of color sums on NVidia GPUs using the ihel3 codebase are compared in Fig. 4, where they are referred to as “ihel3b” (dashed-red) and “ihel3” (solid-red), the only difference between them being an environment variable set at runtime. The picture clearly shows that the BLAS implementation in the ihel3 codebase performs much worse than CUDA kernels for the simpler $g\!\rightarrow\!t\bar{t}, gg$$g\!\rightarrow\!t\bar{t}g$ and $gg\!\rightarrow\!t\bar{t}gg$ physics processes. For the more complex $gg\!\rightarrow\!t\bar{t}ggg$ process, the situation is less clear-cut: for small grids, the kernel implementation is faster for all floating point precisions; for large grids, however, the ihel3 BLAS implementation eventually becomes as fast as the kernel implementation, and in double and float precision (but not in mixed precision) it is eventually faster for very large grids. This is interesting, but of not much practical relevance, as most production workflows would use small grids to keep event generation jobs more manageable. This observation, however, was the starting point of my further refinement of BLAS color sums in the “ihel3p1” version of the code, described in a later subsection. On AMD GPUs, Fig. 5 shows that the kernel version of color sums is always much better than the corresponding hipBLAS implementation, in the ihel3 codebase.
Figure 8: Time spent in the color sum as a fraction of the total time spent in the sigmaKin ME engine for the $gg\!\rightarrow\!t\bar{t}ggg$ process, as a function of the number of events nevt processed in parallel in one GPU iteration. Lower is better. In each plot, four curves are shown, using the ihel3 (red) and ihel3p1 (black) codebases, in the CUDA kernel (solid) and CUDA/BLAS (dashed) implementations. The 3 plots correspond to 3 floating point precisions. The handling of the nGoodHel good helicities is different in the various scenarios. In the two CUDA kernel implementations, nGoodHel kernels are launched in separate helicity streams, each one on a GPU grid of size nevt: in particular, nevt is the product of the number of threads per block, which is fixed to 32, by the varying number of blocks in the GPU grid. For the CUDA/BLAS implementations: in ihel3b, nGoodHel separate BLAS color matrix multiplications are computed, in different helicity streams, using dual amplitude vectors with nevt events; in ihel3b, a single BLAS color matrix multiplications is computed, in the default stream, using dual amplitude vectors with nGoodHel*nevt events. Measurements performed for an NVidia V100 GPU, on the same node used to prepare Fig. 7.
3.3.1 Color sum profiling
To put this work on color sums into context, and to better understand the relative merit of the cuBLAS and kernel implementations in $gg\!\rightarrow\!t\bar{t}ggg$, I found it useful to perform some more detailed profiling of this calculation. One of my aims was to measure the time taken by the color sum as a fraction of the total time taken by the ME calculation in sigmaKin, in different situations. In fact, the motivation for many recent efforts to speed up the color sum calculation, such as our development of the mixed precision mode bib:acat2022 ; bib:chep2024 ; bib:zw2025 or the work I present here on BLAS, is that the profiling bib:kiran of earlier versions of MG5aMC had shown that this could represent up to 60% of the total ME computation for $gg\!\rightarrow\!t\bar{t}ggg$; however, more recent versions of the software, notably CUDACPP, have not yet been profiled in detail.
Initially, I profiled the code by a sampling approach, using perf, but this did not allow detailed color sum profiling on the GPU. I therefore made some additional modifications to the ihel3 version of the code to instrument it with dedicated timers. Specifically, I used some timers based on the x86 rdtsc instruction, which I had developed for some previous profiling work on MG5aMC bib:prof2024 . This approach can provide relatively accurate results with a limited (<<10%) overhead.
The results of my analysis, which are shown in Fig. 7 for the CUDA and SIMD C++ backends, are somewhat surprising: for $gg\!\rightarrow\!t\bar{t}ggg$, in the ihel3 version of the software, the color sum implementation in CUDACPP represents only 5 to 10% of the total time to compute the ME in all SIMD CPU modes, while in the CUDA kernel implementation this fraction is around 28%, 6% and 15% in double, mixed and single precision modes, respectively. These results are particularly good for the mixed precision mode, and indicate that our previous optimizations have managed to reduce the time taken by the color sum to a level where this is no longer a bottleneck of the ME calculation.
In Fig. 7, the results for BLAS color sums are not shown: this is because they heavily depend on the number of events nevt processed in each GPU cycle, which corresponds to the size of the event vector in BLAS and to the GPU grid size for the calculate_jamp kernels. These results are instead shown by the two red curves in Fig. 8, which show that the efficiency of BLAS color sums increases as nevt increases, while that of the kernel implementation is essentially constant. This plot is consistent with the trend shown in Fig. 4 and is useful to better understand it. The two other black curves in Fig. 8 refer to the further improvements in the “ihel3p1” version of the software, described further below.
In the C++ implementation, one reason for the relatively low time footprint of the color sum may be a set of optimizations that I introduced in October 2022 after some discussions in the team (and notably some suggestions by Olivier Mattelaer), and never documented in detail. To start with, as discussed above for BLAS in Eq. 1, I simplified the color sum on the complex vector of dual amplitudes as the sum of two quadratic forms on real vectors. In addition, I also rewrote the multiplication as one involving a triangular matrix, to ensure that the non-diagonal terms of the symmetric matrix are only used once in the calculation, with a factor 2. While writing this paper, I realized that this optimization had been applied to C++ code, but not yet to color sum kernels on GPUs: this is another point that I later fixed in the “ihel3p1” version, described further below.
Color sum implementation Blocks*Threads (nevt) Function FMA Tensor cores
Kernel 8*1 color_sum_kernel 1.9% 0.0%
Kernel 512*32 color_sum_kernel 9.0% 0.0%
cuBLAS 8*1 ampere_sgemm_32x32_sliced1x4_nt 22.7% 0.0%
gemvNSP_kernel<<…cublasGemvTensorStridedBatched…>> 7.3% 0.0%
cuBLAS 512*32 cutlass::Kernel2<<cutlass_80_simt…>> 81.5% 0.0%
gemvx::kernel<<…cublasGemvTensorStridedBatched…>> 20.6% 0.3%
cuBLAS/TF32 8*1 cutlass::Kernel2<<cutlass_80_tensorop…>> 13.7% 16.4%
gemvNSP_kernel<<…cublasGemvTensorStridedBatched…>> 7.3% 0.0%
cuBLAS/TF32 512*32 cutlass::Kernel2<<cutlass_80_tensorop…>> 14.1% 24.4%
gemvx::kernel<<…cublasGemvTensorStridedBatched…>> 20.3% 0.3%
Table 3: Summary of the results for the profiling of the check.exe application, using the NVidia NSight Compute (ncu) profiler version 2024.1.1.0. The results are obtained on an NVidia A100 GPU, using mixed floating precision in the ihel3 version of CUDACPP. Only the kernels in one specific helicity stream were profiled. The columns represent the following: (1) color sum implementation (three configurations are tested: default using CUDA kernels; cuBLAS; cuBLAS with TF32 math mode); (2) GPU grid size configuration for the kernel implementation (or, equivalently, number of events processed in one cuBLAS multiplication); (3) simplified function name printed out by the profiler; (4) FMA (CUDA core) activity, as per ncu metric sm__pipe_fma_cycles_active.avg.pct_of_peak_sustained_active; (5) Tensor core activity, as per ncu metric sm__pipe_tensor_cycles_active.avg.pct_of_peak_sustained_active.
3.3.2 Tensor core profiling
As the use of tensor cores was one of the main motivations for prototyping cuBLAS for color sums, I present in Table 3 the results of some quick studies using the NVidia NSight Compute (ncu) profiler. Unlike the results shown so far in this paper, these were obtained using an NVidia A100 GPU, which has more advanced tensor core features than the older V100 model. The results of this test, which was performed using CUDACPP ihel3 with mixed floating precision, clearly show that tensor cores are used by cuBLAS color sums. In particular, the cublasSgemm multiplication $(C)V$ makes significant use of tensor cores, but only if TF32 math mode is explicitly enabled, as otherwise cuBLAS seems to prefer traditional SIMT kernels. As for the multiplication of $V^{t}$ by $(C)V$ in cublasSgemmStridedBatched, two different implementations are used by cuBLAS for small and large numbers of events, but neither of them makes a significant use of tensor cores. Other tests in different configurations, whose results are not presented here in detail, indicate that cuBLAS color sums make a significant use of tensor cores (irrespective of whether TF32 math mode is enabled or not) when double precision is used. This is somewhat surprising, as tensor cores are designed and optimized for lower precision calculations in AI. A more detailed investigation of tensor cores in the CUDACPP cuBLAS implementation is beyond the scope of this paper, but could be useful for instance in the context of the HEP-SCORE benchmarking project. In particular, further investigations could be useful to understand the usefulness of cuBLAS for CUDACPP color sums on more recent NVidia GPUs, with much more advanced tensor core features.
3.4 Further improvements of color sums (“ihel3p1”)
As mentioned above, while writing the first preprint for this paper bib:preprintV1 I realised that two optimizations of color sums on GPUs could still be attempted. I have implemented both of them in a later version “ihel3p1” of the code, the results for which are represented by the black-solid (ihel3p1, BLAS disabled) and black-dashed lines (ihel3p1b, BLAS enabled) in Figures 4 and 5 for NVidia and AMD GPUs. These patches have already been merged in production with the rest of my first batch of changes to the code: the black lines in the figures, in other words, represent the performance of the current latest release CUDACPP v1.01.01. As shown in Fig. 6, I also note that the C++ implementation in this ihel3p1 release has essentially the same performance as the previous ihel0 release for all SIMD builds.
The first and simplest optimization was that, a previously done in C++, I rewrote the color sum kernel using a triangular matrix, to ensure that the non-diagonal terms of the symmetric matrix are only used once in the calculation with a factor 2, rather than twice. As shown by the bottom-left and bottom-right plots in Fig. 4, this small change does provide a minor but visible increase in peak throughput in $gg\!\rightarrow\!t\bar{t}ggg$, for double and single precision, but not for mixed precision where the time footprint of the color sum was already small. This is also confirmed by comparing the solid-red (ihel3) and solid-black (ihel3p1) curves in Fig. 8, which shows that the fraction of time spent in color sum kernels is reduced by around 5% in both double and single precision.
The second optimization, conversely, concerns BLAS and was motivated by the observations, discussed above and based on Figures 4 and 8, that BLAS throughputs significantly increase when large vectors are used. The basic idea, therefore, was to group together the color sums for all helicities as a single BLAS multiplication, rather than perform a separate BLAS multiplication for each helicity. The changes that I implemented are therfore the following: first, I removed the color sum dispatcher from the helicity loop, implementing a single dispatcher for all helicities rather than one per helicity; in the default case where color sum kernels are used, the dispatcher takes care of launching separate kernels for the different helicities, in the corresponding helicity streams; if BLAS is selected at runtime, however, the dual amplitudes from the different helicities are concatenated to form a single much larger array of events, as if the different helicities for each event were treated as different events, and a single BLAS multiplication is performed in the default stream. This is represented in the lower diagrams of Fig. 11 (this refers to a later ihel6 version of the software, which however only differs from ihel3p1 in the handling of Feynman diagrams). As shown in Fig. 4, the BLAS throughputs of this improved ihel3p1b implementation (dashed-black) are much higher than those of the previous ihel3b implementation (dashed-red). For $gg\!\rightarrow\!t\bar{t}ggg$ (bottom row), in particular, the performance receives a massive boost: BLAS color sums are now better than kernels in double precision, and have equivalent performance in mixed and single precision. This can also be seen in Fig. 8, which shows that the fraction of time spent in the color sum in $gg\!\rightarrow\!t\bar{t}ggg$ is now consistently in the 5–10% range when BLAS is used, independently of the number of events in each GPU iteration. For simpler processes, however, color sum kernels still maintain a higher performance: keeping this into account, it seems appropriate to keep color sum kernels as the default and BLAS as an optional alternative.
Figure 9: Schematic representation of the CUDACPP ME engine (sigmaKin), in the fourth kernel splitting scenario described in this paper: (ihel4) Feynman diagrams as individual kernels.
4 Kernel splitting II: Feynman diagram groups
After implementing helicity-level parallelism, separating Feynman diagrams from color sums and porting the latter to BLAS, the next step in my kernel splitting developments consisted, quite naturally, in trying to further decompose the calculate_jamps kernel, which in ihel3 computes dual amplitudes from all Feynman diagrams, into many shorter kernels. My initial motivation for attempting this was to investigate whether this may lead to further throughput increases, possibly thanks to the lower number of registers used by smaller GPU kernels (lower “register pressure”): I soon realised, however, that I was not able to achieve further throughput increases, because moderately small Feynman diagram kernels are still quite complex and especially come at the cost of a very large increase in GPU global memory access, which massively slows down the calculation. However, there is also a second motivation to splitting Feynman diagrams, namely making the code easier to build and manage for very complex physics processes with very large numbers of diagrams. As described in this section, my work in this direction did achieve some interesting progress, such as the possibility to compute $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes which were previously non-computable in the CUDACPP software.
In the following subsections, I describe my second batch of three sets of kernel splitting changes, which I now also consider ready to be integrated in new production releases of CUDACPP: (ihel4) one kernel per Feynman diagram in each helicity stream; (ihel4p1) ditto, with optional CUDA graph orchestration; (ihel5) one kernel per diagram group in each helicity stream, with separate source code files; (ihel6) a single kernel in each helicity stream, processing all diagram groups as device functions. The results of the intermediate ihel4 step and those of the final ihel6p2 step on $gg\!\rightarrow\!t\bar{t}$ production with 0, 1, 2 or 3 extra final state gluons are given in Figures 4 and 5 for NVidia and AMD GPUs: these plots show, in particular, that the achieved throughputs in the ihel6p2 software are completely equivalent to those of the earlier ihel3p1 software, when the code is generated without diagram splitting, i.e. with a single kernel processing all Feynman diagrams as in ihel3p1. The main interest of ihel6p2, conversely, comes from the more complex $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$ with 4 and 5 extra final state gluons, the results for which are presented at the end of this section.
Figure 10: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU (node itscrd90 at CERN). Higher is better. The 2 plots correspond to throughputs derived by executing the check.exe application through 1 (left) or 10 (right) cycles of a GPU grid. The number of threads per block is fixed to 32 (warp size); the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel3) all Feynman diagrams in a single kernel; (ihel4) Feynman diagrams as individual kernels, without CUDA graphs; (ihel4p1) Feynman diagrams as individual kernels, with CUDA graphs.
4.1 Feynman diagrams as individual kernels (“ihel4”)
The first step in my attempt to split the calculation of all Feynman diagrams in calculate_jamps kernel into smaller GPU kernels consisted, not surprisingly, in encapsulating each diagram in an individual kernel. In practice, rather than launching one calculate_jamps kernel per helicity stream, which internally processes all Feynman diagrams, the new code launches many kernels diagram1, diagram2, …diagramN sequentially in each helicity stream. In both cases, the processing of Feynman diagrams must follow a predefined order, because for every diagram the algorithm knows which intermediate wavefunctions have already been computed, and which wavefunctions must still be computed. This “ihel4” version of the software is illustrated in Fig. 9.
Unfortunately, this version of the software has a consistently worse performance than the previous ones, both on NVidia and AMD GPUs, as shown in Figures 4 and 5. With respect to ihel3, for instance, the ihel4 throughputs for $gg\!\rightarrow\!t\bar{t}ggg$ on an NVidia V100 GPU are a factor 3-4 worse than ihel3p1 at their peak for large grids, and worse by much larger factors for small grids. In addition, ihel4 crashes for very large grids, in configurations where ihel3p1 performs very well. As shown in Fig. 6, the ihel4 software in vectorized C++ for SIMD CPUs also performs worse than ihel3p1, with degradations in throughput up to 40% for $gg\!\rightarrow\!t\bar{t}ggg$.
There are at least two possible explanations for the worse GPU performance of ihel4: heavier access to GPU global memory, and kernel launch overhead. At the time of writing my previous preprint bib:preprintV1 for this article, I wrongly assumed that the latter was the most likely culprit. With hindsight and after many additional tests and developments, it is now clear that the bottleneck is instead the former. Kernel launch overhead can be reduced by adopting the CUDA graphs bib:graphsblog technology: this is what I implemented in the ihel4p1 software, described below. Reducing the impact of GPU global memory access, conversely, required a much more radical rethink of Feynman diagram splitting, described in the two following subsections: first of all, by launching kernels which include groups of many Feynman diagrams (ihel5), rather than individual diagrams; later on, by moving back to a single kernel, which however is no longer monolithic as it internally calls diagram groups as GPU device functions (ihel6).
Figure 11: Schematic representation of the CUDACPP ME engine (sigmaKin), in the last kernel splitting scenario described in this paper: (ihel6) Feynman diagram groups as individual kernels or as device functions in a single kernel.
4.1.1 Small kernels in CUDA graphs (“ihel4p1”)
[…]
The switch to CUDA graphs was relatively easy and was achieved in the ihel4p1 software. As I did for many other components of CUDACPP, I added this feature as an option, which is disabled by default and must be explicitly enabled at runtime if desired [^9] Set environment variable CUDACPP_RUNTIME_GPUGRAPHS at runtime to enable CUDA graphs (Nvidia GPUs) or HIP graphs (AMD GPUs) orchestration of Feynman diagram kernels.. In practice, I did the following: the calculate_jamps host function was restructured so that it delegates kernel launching of Feynman diagrams to a gpuDiagrams wrapper function; calculate_jamps also keeps a static set of pointers that are needed to configure graphs if required; if graphs are disabled (default), gpuDiagrams is a pass-through function that simply launches each kernel directly in the relevant helicity stream; if graphs are enabled, however, gpuDiagrams configures graphs on its first execution (i.e. if the static configuration in calculate_jamps is empty), while on subsequent executions it runs kernels via graphs rather than launching them individually.
The performance results for this implementation can be seen in Fig. 10, by comparing the grey (ihel4, no graphs), cyan (ihel4p1, graphs enabled) and red (ihel3, single kernel) curves, for a single execution (left) or ten executions (right) of the workflow. As expected, these curves indicate that graphs slow down the workflow when kernels are executed only once (ihel4p1 is worse than ihel4 in the left plots), but they speed it up when the workflow is executed several times (ihel4p1 is worse than ihel4 in the right plots). However, the plots also show that, even if graphs are used, launching Feynman diagrams as individual kernels is still much slower than keeping them together in a single kernel (ihel4p1 is much worse than ihel3 in both plots). This is a clear indication that the main problem in this approach is not the kernel launch overhead, but global memory access: this is precisely the issue addressed using diagram groups in the ihel5 and ihel6 versions of the software, which are described in the two next subsections.
4.2 Feynman diagram groups as kernels (“ihel5”)
After showing that the impact of kernel launch overhead was relatively limited, it was thus clear that the most likely explanation for the worse GPU performance of ihel4 with respect to ihel3 is memory access. One technical challenge in this case, in fact, is that the wavefunctions for all external and internal particles, which are a local variable wf in the calculate_wavefunctions and calculate_jamps kernels in versions ihel0–ihel3 of the code (see Table 1), must now also be stored into and retrieved from GPU global memory. In addition, as previously mentioned in a footnote, the existing implementation of helicity amplitude functions like FFV1_0 essentially constrains the wavefunction layout in GPU global memory to be an Array-Of-Structures (AOS) wf[nevt][6][2], which is suboptimal as it does not allow coalesced memory access as an SOA would do. This further increases the pressure on memory bandwidth, and may represent an additional cause of slowdowns.
Figure 12: Throughputs (ME/s) as a function of grid size for $gg\!\rightarrow\!t\bar{t}ggg$ on an NVidia V100 GPU. Higher is better. The number of threads per block is fixed to 32 (warp size); the grid size is varied by changing the number of blocks. Each plot compares code generated with different numbers of diagrams per group (dpg), ranging from 1 to 2000. The thick solid line (dpg2000) corresponds to the case where all 1240 diagrams are in a single diagram group, i.e. to the default scenario for $gg\!\rightarrow\!t\bar{t}ggg$ in the rest of this paper. The 2 plots correspond to code built in the DCDIAG=0 (left: each diagram group is a kernel) and DCDIAG=1 (right: there is a single kernel, calling all diagram groups as device functions) build modes.
The next step in my kernel splitting developments, as a consequence, was an attempt to reduce the frequency of GPU global memory access, while still partitioning the full list of Feynman diagrams into several smaller and more manageable kernels. My “ihel5” implementation, in particular, consisted in grouping together many Feynman diagrams in kernels of intermediate sizes. In doing this, I kept unchanged the order in which Feynman diagrams are sequentially computed. As I did not reorganise the calculation to allow some level of parallelism, I did not expect that these developments could yield any throughput increases. Their interest, instead, was mainly that they could make the software more modular and manageable, possibly leading to faster build times and possibly allowing ME calculations with a large number of final state particles. This goal was indeed achieved, as detailed below.
My ihel5 implementation of diagram groups is configurable according to two parameters: one “dpg” parameter defines the maximum number of Feynman diagrams in one diagram group, while another “dpf” parameter defines the minimum number of Feynman diagrams in one source code file. Currently these parameters are only managed through special scripts, but eventually they could become standard MG5aMC runcard parameters. In practice, diagram group kernels are implemented as follows: for each diagram group, the code generator keeps track of which wavefunctions the diagram group needs in its calculations, and of which wavefunctions the diagram group updates; at the beginning of all diagram groups except the first one, those wavefunctions are read from global memory and copied to local variables in GPU registers; incidentally, this makes it possible to define a better SOA layout in global memory for coalesced memory access, while maintaining the legacy AOS layout wf[nevt][6][2] in local variables to avoid the need to restructure all calculations; at the end of all diagram groups except the last one, any updated wavefunctions are written back to GPU global memory; at the end of all diagram groups, the jamps dual amplitudes in GPU global memory are also updated from the results of those computed using local variables. This design implies that, the lower the number of diagram groups, the lower the frequency of GPU global memory access. In particular, in the special case where the dpg is larger than the total number of diagrams in the physics process, only one diagram group is created: this is executed as a single kernel, which does not need to read wavefunctions at the beginning or to write them back at the end, so effectively this is a single-kernel implementation equivalent to that in ihel3, and indeed has the same performance, as previously noted.
For reasons clarified below, ihel5 is equivalent to the default DCDIAG=0 implementation in the ihel6 version of the software, which is illustrated in the upper-left part of Fig. 11. As a consequence, in this paper I do not show explicit results for the ihel5 software, but only those from ihel6. The impact of the dpg parameter on ihel5 performance for $gg\!\rightarrow\!t\bar{t}ggg$, in particular, is shown by the left plot in Fig. 12, derived from the default DCDIAG=0 version of ihel6. The plot is interesting as it shows that peak performance, equivalent to that of ihel3, can only be achieved with a single kernel that processes all diagrams: even a minimal splitting of the 1240 diagrams into two diagram groups (dpg=1000), in fact, comes at the cost of a very large performance hit for small GPU grids and of a 10% reduction in peak performance for large grids (with the additional problem that the program crashes due to limited memory when the GPU grids are too large). This clearly shows that GPU global memory access is very expensive.
4.3 Diagram groups as device functions (“ihel6”)
Having understood that processing all Feynman diagrams in a single kernel is better for performance, but still wishing to split up complex physics processes into smaller and more manageable software chunks, my next step of development was a relatively minor, but rather consequential, modification of ihel5 diagram groups. As I had already prepared the code generator to define diagram groups, I simply transformed diagram groups into GPU device functions and turned calculate_jamps into a single GPU kernel that calls them all sequentially. In this “ihel6” software version, illustrated in the upper-right part of Fig. 11, the big difference is that all wavefunctions are defined as local variables of GPU device functions, stored in GPU registers, and do not need to be read from or stored into GPU global memory. Sticking to my general strategy of allowing the choice between different alternatives, the code generated in ihel6 supports both the ihel5-like mode, where diagram groups are GPU kernels, and the new ihel6-like mode, where diagram groups are device functions. The choice between the two implementations can be made at build time, by passing DCDIAG=1 to specify that device code diagrams should be used, or by keeping the default value DCDIAG=0. The name of this variable is also meant as a reminder that, for technical reasons related to the handling of coupling constants and other physics parameters, GPU relocatable device code (i.e. the CUDA rdc option) is needed in the DCDIAG=1 implementation.
The impact on performance of the DCDIAG=1 approach is quite interesting, as shown in the right plot in Fig. 12. As in the case of DCDIAG=0, performance is quite poor for low values of dpg, i.e. when there are many diagram groups. Unlike DCDIAG=0, however, peak performance, only slightly lower than that of ihel3, is reached for intermediate values of dpg between 100 and 1000. For dpg=2000, instead, i.e. with a single diagram group, performance degrades again. In other words, there seems to be a sweet spot for diagram groups of intermediate sizes, where DCDIAG=1 performance is reasonably good. This is especially interesting because this is the same sweet spot where source code files and their builds appear to be easier to manage. This is further discussed in the following, in the case of the more complex processes $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$.
$gg\!\rightarrow\!t\bar{t}gggg$ with 15495 diagram groups (1 dpg, 100 dpf) — build takes 2m20s
C++ SIMD build mode none sse4 avx2 512y 512z
Time to compute 16 MEs 33.9s 37.8s 32.8s 31.6s 30.3s
Throughput 0.47/s 0.42/s 0.49/s 0.51/s 0.53/s
Speedup w.r.t. none x1 x0.90 x1.03 x1.07 x1.12
$gg\!\rightarrow\!t\bar{t}gggg$ with 1550 diagram groups (10 dpg, 100 dpf) — build takes 3m10s
C++ SIMD build mode none sse4 avx2 512y 512z
Time to compute 16 MEs 9.57s 10.27s 4.59s 4.89s 3.83s
Throughput 1.67/s 1.56/s 3.48/s 3.27/s 4.18/s
Speedup w.r.t. none x1 x0.93 x2.08 x1.96 x2.50
$gg\!\rightarrow\!t\bar{t}gggg$ with 155 diagram groups (100 dpg, 100 dpf) — build takes 3m
C++ SIMD mode none sse4 avx2 512y 512z
Time to compute 16 MEs 7.18s 4.52s 1.73s 1.53s 1.05s
Throughput 2.23/s 3.53/s 9.22/s 10.0/s 15.2/s
Speedup w.r.t. none x1 x1.59 x4.14 x4.69 x6.83
$gg\!\rightarrow\!t\bar{t}gggg$ with 78 diagram groups (200 dpg, 200 dpf) — build takes 3m10s
C++ SIMD mode none sse4 avx2 512y 512z
Time to compute 16 MEs 6.30s 3.46s 1.45s 1.40s 0.89s
Throughput 2.54/s 4.63/s 11.0/s 11.4/s 18.0/s
Speedup w.r.t. none x1 x1.82 x4.33 x4.50 x7.08
$gg\!\rightarrow\!t\bar{t}gggg$ with 16 diagram groups (1000 dpg, 1000 dpf) — build takes 3m40s
C++ SIMD build mode none sse4 avx2 512y 512z
Time to compute 16 MEs 6.23s 3.08s 1.29s 1.22s 0.72s
Throughput 2.57/s 5.19/s 12.4/s 13.1/s 22.2/s
Speedup w.r.t. none x1 x2.02 x4.81 x5.11 x8.65
$gg\!\rightarrow\!t\bar{t}gggg$ with 2 diagram groups (10000 dpg, 10000 dpf) — build takes 28m
C++ SIMD build mode none sse4 avx2 512y 512z
Time to compute 16 MEs 6.05s 2.93s 1.22s 1.17s 0.70s
Throughput 2.64/s 5.45/s 13.1/s 13.7/s 22.9/s
Speedup w.r.t. none x1 x2.06 x4.97 x5.18 x8.66
$gg\!\rightarrow\!t\bar{t}gggg$ with 1 diagram group (100000 dpg, 100000 dpf) — build FAILS
g++: internal compiler error: Segmentation fault
Please submit a full bug report, with preprocessed source if appropriate
Table 4: C++ run time performance of $gg\!\rightarrow\!t\bar{t}gggg$ in different SIMD modes and with different numbers of diagrams per group (FPTYPE=m). C++ code generated using CUDACPP hack_ihel6p2. Results for a single core of an Intel Xeon Gold 6326 CPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering, which is still performed sequentially one helicity at a time. As a basic validation of physics results, it was only checked that all calculations return exactly the same average ME value, when starting from the same random number seeds, for all SIMD and diagram-per-group modes.
$gg\!\rightarrow\!t\bar{t}gggg$ with 15495 diagram groups (1 dpg, 100 dpf)
Build mode Build time Throughputs in MEs/s (blocks * threads)
1*32 2*32 4*32 8*32 16*32 32*32 64*32 128*32
DCDIAG=0 48m 3.2E-2 6.4E-2 1.3E-1 2.4E-1 4.0E-1 fail fail fail
DCDIAG=1 10m 6.5E-1 1.2E0 2.1E0 3.1E0 4.2E0 4.9E0 5.4E0 5.7E0
$gg\!\rightarrow\!t\bar{t}gggg$ with 1550 diagram groups (10 dpg, 100 dpf)
Build mode Build time Throughputs in MEs/s (blocks * threads)
1*32 2*32 4*32 8*32 16*32 32*32 64*32 128*32
DCDIAG=0 4m30s 6.6E-1 1.4E0 2.6E0 3.2E0 4.8E0 fail fail fail
DCDIAG=1 3m50s 1.7E1 2.8E1 4.2E1 4.7E1 5.1E1 5.1E1 5.2E1 5.2E1
$gg\!\rightarrow\!t\bar{t}gggg$ with 155 diagram groups (100 dpg, 100 dpf)
Build mode Build time Throughputs in MEs/s (blocks * threads)
1*32 2*32 4*32 8*32 16*32 32*32 64*32 128*32
DCDIAG=0 8m50s 2.9E0 5.9E0 1.1E1 2.1E1 4.0E1 fail fail fail
DCDIAG=1 6m20s 7.0E1 1.1E2 1.7E2 2.1E2 2.4E2 2.4E2 2.4E2 2.4E2
$gg\!\rightarrow\!t\bar{t}gggg$ with 78 diagram groups (200 dpg, 200 dpf)
Build mode Build time Throughputs in MEs/s (blocks * threads)
1*32 2*32 4*32 8*32 16*32 32*32 64*32 128*32
DCDIAG=0 15m 4.1E0 9.3E1 1.8E1 3.5E1 6.6E1 fail fail fail
DCDIAG=1 17m fail fail fail fail fail fail fail fail
$gg\!\rightarrow\!t\bar{t}gggg$ with 16 diagram groups (1000 dpg, 1000 dpf)
Build mode Build time Throughputs in MEs/s (blocks * threads)
1*32 2*32 4*32 8*32 16*32 32*32 64*32 128*32
DCDIAG=0 2h57m 6.5E0 1.2E1 2.4E1 4.7E1 8.9E1 fail fail fail
DCDIAG=1 failed (killed by oom-kill) N/A N/A N/A N/A N/A N/A N/A N/A
Table 5: CUDA run time performance of $gg\!\rightarrow\!t\bar{t}gggg$ in different DCDIAG modes and with different numbers of diagrams per group (FPTYPE=m). CUDA/C++ code generated using CUDACPP ihel6p2. Results for an NVidia A100 GPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering, which is still performed sequentially one helicity at a time. As a basic validation of physics results, it was only checked that all calculations return exactly the same average ME value, when starting from the same random number seeds, for all SIMD and diagram-per-group modes.
Before moving on to more complex physics processes that can now be computed by splitting the calculation into several diagram groups, I add one final comment about our standard candles, $g\!\rightarrow\!t\bar{t}, gg$$\rightarrow\!t\bar{t}g, gg$$g\!\rightarrow\!t\bar{t}gg$ and $gg\!\rightarrow\!t\bar{t}ggg$. In particular, as shown in Figures 4 and 5 for NVidia and AMD GPUs and in Fig. 6 for an Intel CPU, I stress that the ihel6 line of software versions, and more specifically the “ihel6p2” version described below, performs just as well as the current ihel3p1 release, or even slightly better, when these processes are generated using a single diagram group. In other words, the new diagram splitting functionality in ihel6p2 offers the benefit of allowing computations of more complex processes, without compromising the well established performance for simpler processes.
4.4 Studies of a $2\!\rightarrow\!6$ process: $gg\!\rightarrow\!t\bar{t}gggg$
Using the new features in CUDACPP ihel6, notably its diagram splitting capabilities and the possibility to treat diagram groups either as independent kernels or as device methods on GPUs, I was able to test and obtain some throughput results for complex $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ physics processes such as $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$. I note that some studies of $gg\!\rightarrow\!t\bar{t}gggg$ had already been performed in the past using the previous ihel0 software with a monolithic sigmaKin kernel, by myself and other members of the development team, but they had immediately appeared extremely challenging. In my recollection, in 2023 I had been able to build the C++ code for $gg\!\rightarrow\!t\bar{t}gggg$, but only using the clang compiler, and this had taken 32 hours. The C++ build using g++, conversely, had crashed with an internal error, while the CUDA build had gone on for more than one week before I decided to stop it.
Diagram splitting dramatically improves this situation, not only for the GPU implementation that I was initially targeting, but also for the vectorised C++ implementation. As shown in Table 4, the C++ build of $gg\!\rightarrow\!t\bar{t}gggg$ with g++ successfully completes in less than 4 minutes if the number dpg of diagrams per group is kept below 1000, and it takes longer but still completes in 30 minutes if dpg is increased to 10000, which implies only two diagram groups for this specific process. Interestingly, instead, the g++ build fails with an internal compiler error if all 15495 diagrams are included in the same function: this is exactly the same situation that I had faced earlier on with the ihel0 software. Table 4 is also very interesting because it shows that C++ runtime throughputs are significantly degraded with fewer than 100 dpg, but they are essentially at the their peak plateau with 1000 dpg: the runtime performance with 10k dpg is only negligibly better than that with 1k dpg, at the cost of a much longer build time. This observation, which seems to fall in line with that previously made about $gg\!\rightarrow\!t\bar{t}ggg$ for GPUs, is especially interesting because it seems to suggest that, if the C++ build of the 100k dpg code (i.e. that with a monolithic function) had succeeded, its runtime performance might have been only negligibly better than that with 10k or 1k dpg. In other words, also for the C++ implementation of $gg\!\rightarrow\!t\bar{t}gggg$ (similarly to what was previously observed for the GPU implementation of $gg\!\rightarrow\!t\bar{t}ggg$), there seems to be a sweet spot around 1k dpg where software builds are fast and manageable, and runtime performance is satisfactory. It should be noted, in particular, that the SIMD speedup observed with dpg 1k or 10k is consistent with the maximum achievable in this configuration, using mixed floating point precision: on this Intel Xeon Gold 6326 CPU with two FMA units, one would expect a maximum SIMD speedup of x8 for Feynman diagrams and x16 for the color matrix, and an overall speedup somewhere in between, which is consistent with the observed speedups around x9. This very good scaling seems to suggest that this calculation is not limited by memory access on the CPU.
$gg\!\rightarrow\!t\bar{t}gggg$ with 155 diagram groups (100 dpg)
C++ SIMD build mode none sse4 avx2 512y 512z
Time to compute 32 MEs 13.8s 8.91s 3.26s 3.21s 2.18s
$\rightarrow$ Time in calculate_jamps 12.5s (90.1%) 8.24s (92.6%) 2.95s (90.3%) 2.89s (90.2%) 2.01s (92.2%)
$\rightarrow$ Time in color_sum_cpu 1.36s (9.8%) 0.66s (7.4%) 0.32s (9.7%) 0.32s (9.8%) 0.17s (7.8%)
Throughput 2.31/s 3.59/s 9.81/s 9.97/s 14.7/s
Mean ME value / E-10 3.057124 3.057124 3.057124 3.057124 3.057124
$gg\!\rightarrow\!t\bar{t}gggg$ with 16 diagram groups (1000 dpg)
C++ SIMD build mode none sse4 avx2 512y 512z
Time to compute 32 MEs 13.3s 6.08s 2.55s 2.46s 1.43s
$\rightarrow$ Time in calculate_jamps 12.0s (89.8%) 5.46s (89.9%) 2.23s (87.7%) 2.14s (87.2%) 1.27s (88.7%)
$\rightarrow$ Time in color_sum_cpu 1.36s (10.2%) 0.62s (10.1%) 0.31s (12.2%) 0.31s (12.7%) 0.16s (11.3%)
Throughput 2.40/s 5.27/s 12.6/s 13.0/s 22.4/s
Mean ME value / E-10 3.057124 3.057124 3.057124 3.057124 3.057124
Table 6: Profiling of Feynman diagrams and color sums for $gg\!\rightarrow\!t\bar{t}gggg$ on CPU in different SIMD modes and with different numbers of diagrams per group (FPTYPE=m). C++ code generated using a modified version of CUDACPP, instrumented with rdtsc timers (hack_ihel6p2_ggtt4g branch). Results for a single core of an Intel Xeon Gold 6326 CPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering. The average ME value shown in the table, computed from the same random numbers in all cases, is used as a basic validation of physics results.
$gg\!\rightarrow\!t\bar{t}gggg$ with 1000 dpg — (1) diagrams: DCDIAG=0 — (2) color sum: kernel
MEs per cycle (blocks * threads) 32 (1*32) 128 (4*32) 512 (16*32)
Time to compute one ME cycle 5.16s 5.20s 5.66s
$\rightarrow$ Time in calculate_jamps 5.02s (97.3%) 5.09s (95.8%) 5.25s (92.7%)
$\rightarrow$ Time in color_sum_gpu 0.14s (2.7%) 0.22s (4.2%) 0.41s (7.3%)
Throughput 6.20/s 24.1/s 90.5/s
Mean ME value / E-10 3.057124 5.683925 264.1109
$gg\!\rightarrow\!t\bar{t}gggg$ with 100 dpg — (1) diagrams: DCDIAG=0 — (2) color sum: kernel
MEs per cycle (blocks * threads) 32 (1*32) 128 (4*32) 512 (16*32)
Time to compute one ME cycle 9.54s 11.8s 12.4s
$\rightarrow$ Time in calculate_jamps 9.40s (98.6%) 11.6s (98.1%) 12.0s (96.7%)
$\rightarrow$ Time in color_sum_gpu 0.14s (1.4%) 0.22s (1.9%) 0.42s (3.3%)
Throughput 3.35/s 10.9/s –/s
Mean ME value / E-10 3.057124 5.683925 264.1109
$gg\!\rightarrow\!t\bar{t}gggg$ with 100 dpg — (1) diagrams: DCDIAG=1 — (2) color sum: kernel
MEs per cycle (blocks * threads) 32 (1*32) 128 (4*32) 512 (16*32)
Time to compute one ME cycle 0.48s 0.76s 2.14s
$\rightarrow$ Time in calculate_jamps 0.32s (68.1%) 0.50s (66.1%) 1.71s (80.1%)
$\rightarrow$ Time in color_sum_gpu 0.15s (31.8%) 0.26s (33.9%) 0.43s (19.9%)
Throughput 66.1/s 167.6/s 239.7/s
Mean ME value / E-10 3.057124 5.683925 264.1109
$gg\!\rightarrow\!t\bar{t}gggg$ with 100 dpg — (1) diagrams: DCDIAG=1 — (2) color sum: cuBLAS
MEs per cycle (blocks * threads) 32 (1*32) 128 (4*32) 512 (16*32)
Time to compute one ME cycle 0.34s 0.54s 1.76s
$\rightarrow$ Time in calculate_jamps 0.31s (90.9%) 0.50s (93.9%) 1.71s (97.0%)
$\rightarrow$ Time in color_sum_gpu 0.031s (9.1%) 0.032s (6.1%) 0.054s (3.0%)
Throughput 92.8/s 238.7/s 290.6/s
Mean ME value / E-10 3.057124 5.683925 264.1109
$gg\!\rightarrow\!t\bar{t}gggg$ with 100 dpg — (1) diagrams: DCDIAG=1 — (2) color sum: cuBLAS/TF32
MEs per cycle (blocks * threads) 32 (1*32) 128 (4*32) 512 (16*32)
Time to compute one ME cycle 0.35s 0.54s 1.75s
$\rightarrow$ Time in calculate_jamps 0.32s (91.4%) 0.51s (94.3%) 1.71s (97.7%)
$\rightarrow$ Time in color_sum_gpu 0.030s (8.5%) 0.031s (5.7%) 0.039s (2.3%)
Throughput 91.5/s 237.1/s 293.4/s
Mean ME value / E-10 3.057422 5.684374 264.1716
Table 7: Profiling of Feynman diagrams and color sums for $gg\!\rightarrow\!t\bar{t}gggg$ on GPU in different DCDIAG modes and with different numbers of diagrams per group. CUDA/C++ code generated using a modified version of CUDACPP, instrumented with rdtsc timers (hack_ihel6p2_ggtt4g branch). Results for an NVidia A100 GPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering. The average ME value shown in the table, computed from the same random numbers in all cases, is used as a basic validation of physics results.
For what concerns the GPU implementation, diagram splitting is also a game changer, allowing the first (to my knowledge) CUDACPP calculation of matrix elements on GPUs for $gg\!\rightarrow\!t\bar{t}gggg$. As shown in Table 5, CUDA builds succesfully complete in less than 10 minutes, in both modes DCDIAG=0 and DCDIAG=1, if the number of diagrams per group is kept below 100. If dpg is increased to 1000, the DCDIAG=0 build succeeds in 3h, while the DCDIAG=1 build fails because it requires too much memory: it was specifically this point that motivated me to get rid of templates, in an attempt to simplify the task of the CUDA compiler, but unfortunately the situation does not improve after removing templates (ihel6p2). This is unfortunate, because the DCDIAG=1 build clearly gives higher peak throughputs, and reaches them with smaller grid sizes, compared to DCDIAG=0. This is not surprising, because the latter involves much heavier read/write access of wavefunctions and dual amplitudes in GPU global memory, while the former only uses GPU registers in most cases.
4.4.1 Profiling of color sums in $gg\!\rightarrow\!t\bar{t}gggg$
The results I have presented for $gg\!\rightarrow\!t\bar{t}gggg$ in Tables 4 and 5 refer to the overall throughput for computing MEs, which includes the calculations of both Feynman diagrams and QCD color sums. Since kernel splitting into diagram groups and the two different DCDIAG options are only relevant to the former and not the latter, I found it useful to separately profile these two components. In doing this, I also tested the impact of the various color sum options in $gg\!\rightarrow\!t\bar{t}gggg$. I only studied a limited number of scenarios, including in particular those which had achieved the highest overall throughputs in the tests presented above.
The results for the SIMD implementation, shown in Table 6, show that the choice of dpg only affects the calculation of Feynman diagrams, as expected. The fraction of time spent in color sums is around 10–12% in the dpg scenario with the fastest calculation of Feynman diagrams: this number is somewhat lower than I had expected, but it is still significantly larger than the 5–10% observed for the simpler $gg\!\rightarrow\!t\bar{t}ggg$ process. The SIMD speedups of the color sum alone, which is computed here in single precision (FPTYPE=m), are also lower than I had expected: for instance, a factor around 9 is observed for the 512z implementation, which is closer to the x8 speedup expected in double precision than to the x16 speedup expected in single precision. This may deserve further investigations in the future [^10] From a first quick analysis, the problem may be that, for FPTYPE=m on CPUs, the quadratic form $J^{H}(C)J$ is maybe not fully computed in single precision: the first multiplication $H(C)J$ is computed in single precision, but the second multiplication $J^{H}\left[(C)J\right]$ seems to be computed in double precision. .
The results for the GPU implementation, listed in Table 7, also show that the large throughput variation previously noted in Table 5 for different choices of dpg and DCDIAG only come from the calculation of Feynman diagrams. In the fastest scenario (100 dpg and DCDIAG=1), in particular, color sums for relatively large grids with 512 events represent 20% of the time needed for the ME calculation in their default implementation with GPU kernels. What is very interesting, however, is that using cuBLAS in this case speeds up color sums by almost one order of magnitude, reducing them to only 3% of the total time taken to compute MEs. This provides additional evidence that cuBLAS may be the best choice for the color sums of particularly complex processes, even if it should not be used as the default as its performance for simple processes is very poor.
In this context, it should also be noted that cuBLAS color sums provide physics results with the same accuracy as those of GPU kernels, except in TF32 math mode. In Tables 6 and 7, in particular, I explicitly included the average ME value of my simple tests: this is the same to seven significant digits in all scenarios I considered both on CPUs and GPUs, when starting from the same random numbers; the only exception is cuBLAS/TF32, where discrepancies of order 10-4 are observed. As discussed in Sec. 3.3, the TF32 option is a switch that I introduced to encourage cuBLAS to use tensor cores. As shown in 7, this does provide a moderate speedup of color sums by approximately 30%, but it seems quite clear that the overall benefit (reducing the color sum fraction from 3% to 2.3%) is not worth the cost of a loss in physics precision.
$gg\!\rightarrow\!t\bar{t}ggggg$ (1000 diagrams per group and per file) — build takes 55m
C++ SIMD mode none sse4 avx2 512y 512z
Time to compute 16 MEs 472.7s 337.3s 214.9s 176.2s 179.8s
$\rightarrow$ Time in calculate_jamps 397.9s (84.2%) 305.6s (90.6%) 198.6s (92.4%) 160.4s (91.1%) 171.2s (95.2%)
$\rightarrow$ Time in color_sum_cpu 74.8s (15.8%) 31.7s (9.4%) 16.3s (7.6%) 15.7s (8.9%) 8.7s (4.8%)
Throughput 3.38E-2/s 4.74E-2/s 7.45E-2/s 9.08E-2/s 8.90E-2/s
Speedup w.r.t. none x1 x1.40 x2.20 x2.68 x2.63
Mean ME value / E-13 1.748038 1.748038 1.748038 1.748038 1.748038
Table 8: Run time performance, including profiling of Feynman diagrams and color sums, of $gg\!\rightarrow\!t\bar{t}ggggg$ in different SIMD modes (FPTYPE=m). C++ code generated using CUDACPP hack_ihel6p1, with 1000 diagrams per group and per file (this is the $gg\!\rightarrow\!t\bar{t}ggggg$ code described in Table 2). The code was later instrumented with rdtsc timers for color sum profiling. Results for a single core of an Intel Xeon Gold 6326 CPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering, which is still performed sequentially one helicity at a time. As a basic validation of physics results, it was only checked that the five calculations return exactly the same average ME value, when starting from the same random number seeds.
4.5 Studies of a $2\!\rightarrow\!7$ process: $gg\!\rightarrow\!t\bar{t}ggggg$
Taking this one step further, using the diagram splitting in ihel6p1 I also performed a first quick analysis of the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$. Since code generation for this process takes 30h, I only attempted one configuration that I thought could be promising, using 1000 diagrams per group. It turns out that, both for the C++ and CUDA builds, the main challenge in this case came not from Feynman diagrams but from the size of the 5k x 5k color matrix, which in CUDACPP is defined as a constexpr. In the case of C++, I was able to address the issue by modifying the constexpr limit in the compiler: this was enough for the build to succeed (in less than 1h), and I was then able to obtain the first CUDACPP results from C++ for a $2\!\rightarrow\!7$ process, which are shown in Tab. 8. In the case of CUDA, I used a similar trick to successfully complete compilation (replacing constexpr compile-time evaluations by runtime const evaluations), but linking failed with the relocation truncated to fit error, which signals that the executable would be too large for the linker to handle. I tried several options to bypass this issue, but none of them succeeded. Concerning runtime performance of the C++ code, the results in Tab. 8 are somewhat puzzling, as the observed SIMD speedups are much lower than one would expect. Unlike the $gg\!\rightarrow\!t\bar{t}gggg$ calculation, this seems to suggest that CPU memory access may be a limiting factor for this calculation. In the table, I also provide the detailed profiling of the Feynamn diagram and color sum components, derived with the same tools used for simpler processes: the fact that color sums have a much better SIMD speedup than Feynman diagrams, and that their overall footprint is relatively low (around 5 to 15%) is an interesting and somewhat surprising indication that the Feynman diagram calculation is the bottleneck for such a complex calculation.
Finally, I should note that this might well be regarded as an academic exercise. To achieve the calculation of processes with a large number of final state gluons, I believe that Berends-Giele recursion relations bib:bg represent a better approach to explicit Feynman diagrams, which has already been successfully ported to GPUs by different generator teams bib:bg-bothmann ; bib:pepper ; bib:martinez . I still find it interesting, in any case, to explore how far one can push calculations based on Feynman diagrams, to identify some practical bottlenecks in the software and hardware we use, but also because Feynman diagram calculations may provide a complementary tool to cross-check other approaches. Diagram splitting, in particular, appears to be a very promising approach to achieve more complex calculations using vectorized C++ code.
4.6 Discussion
To conclude, taking into account that the ihel6p2 software has the same performance without diagram splitting as the latest CUDACPP release based on ihel3p1 for our standard candles such as $gg\!\rightarrow\!t\bar{t}ggg$, but that it also allows the calculation of more complex physics processes such as $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$ when diagram splitting is enabled, my recommendation is that this work should also be merged upstream and integrated in a new CUDACPP release. I believe that the new features offer many opportunities for further studies, certainly from a software point of view but also from a physics perspective. A few technical details would still need to be sorted out, such as the visibility of the dpg, dpf and DCDIAG parameters through standard runcards, or the preparation of new CI tests including diagram splitting. To ease future maintenance, one may also choose to exclude some of the features I developed: in particular, I think that testing CUDA graphs has been a very useful exercise, but the benefits of this technology for Feynman diagrams appear quite limited, and it might therefore be appropriate to remove this specific feature before merging this work upstream.
In addition, in case of stronger interest in the calculation of more complex processes. one issue that may deserve some minor additional development is the following: currently, helicity streams are used only for computing MEs after helicity filtering, but helicity filtering itself, i.e. the selection of the good helicities with non-zero ME contributions, is done sequentially for the various helicities. As a consequence, helicity filtering is extremely long for the $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes I studied, and for this reason I added a hack [^11] Set CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering. The ME calculation will then use all helicities. to bypass helicity filtering (this made sense for $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$ as all helicities do contribute to the ME calculation in that case). A simple solution, which would be very simple to develop but I skipped for lack of time, would be to use parallel helicity streams also during the helicity filtering step. An even better approach might be to study how many streams are actually useful, and limit them to a fixed number, properly chosen taking into account the GPU hardware specs.
5 Summary and outlook
In summary, in this paper I have described my work on two large batches of optimizations of the CUDACPP plugin for the Madgraph5_aMC@NLO (MG5aMC) generator on GPUs and vector CPUs. The new approach mainly consists in splitting the calculation of matrix elements (ME), which had been so far performed using a single large GPU kernel, into smaller kernels. The software I developed and the studies I presented have all been been performed using the CUDACPP standalone application, but are expected to provide benefits also and especially in the full MadEvent-based MG5aMC generator workflow. This work has mainly targeted NVidia GPUs, but benefits the CUDACPP implementation for CPUs, too. It has also been ported to AMD GPUs, but without any attempt to understand or improve its performance on this platform. I have also taken the opportunity of this paper to describe more in detail some features of the CUDACPP software that had not yet been documented, especially those most relevant to these new developments.
The first large batch of changes, which includes three sets of optimizations, has already been merged upstream and integrated in a new production release of the CUDACPP software bib:cudacpp364 , in collaboration with my colleagues. The first optimization, which parallelizes the calculation for different helicities of the initial and final state particles to different CUDA Streams, achieves a reduction by one to two orders of magnitude in the number of events that must be computed in parallel to make an efficient use of the GPU. This is interesting from a user perspective, as it should allow event generation jobs on GPUs using smaller numbers of events than the current CUDACPP. The second and third optimizations reorganize the calculation of color sums as a separate GPU kernel, which is taken as the default implementation, and also offer the possibly to perform it using GPU-specific BLAS libraries. On NVidia GPUs, in particular, the cuBLAS implementation has been optimized to a level where it is faster than kernels for moderately complex processes such as $gg\!\rightarrow\!t\bar{t}ggg$; this is also interesting as it may offer a mechanism for CUDACPP to exploit tensor cores on NVidia GPUs, which are otherwise unused.
The second large batch of enhancements, which includes three further sets of software changes, was still work in progress at the time of my previous preprint bib:preprintV1 of this paper, but it has now been completed and I also recommend its inclusion in a new production release of the software. It is available as a github tag bib:ihel6tag , and a pull request for its merge upstream has been created. These three sets of changes all consist in different ways to split up the computation of Feynman diagrams into separate GPU kernels. The fourth development, which uses a different GPU kernel for each Feynman diagram, has a very poor performance, even when the workflow is orchestrated by CUDA graphs, because it implies very heavy access to GPU global memory. The fifth development allows the definition of kernels that compute larger groups of Feynman diagrams, possibly defined in separate source code files. The sixth development goes back to a single GPU kernel for all Feynman diagrams, which however delegates their calculation to different device functions for separate diagram groups. All these changes are fully functional and have been extensively tested on all platforms.
While this second line of R&D started off as an attempt to further improve throughputs for our standard candles such as $gg\!\rightarrow\!t\bar{t}ggg$, it now seems clear to me that these processes are best computed using the traditional CUDACPP approach of a single GPU kernel for all Feynman diagrams, since they are simple enough to afford this: the same approach and the same throughputs are still achievable using the new software, which can be configured to produce a single diagram group. The interest of the new feature, conversely, is in much more complex physics calculations such as those required by $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes. I have shown, in particular, that diagram splitting makes it possible to compute matrix elements for $gg\!\rightarrow\!t\bar{t}gggg$ in both the GPU and vector CPU implementations of CUDACPP, and for $gg\!\rightarrow\!t\bar{t}ggggg$ in the latter. The code I used for the $gg\!\rightarrow\!t\bar{t}gggg$ tests that I presented is also available as a github tag bib:ihel6tag4g , which I created only for documentation purposes and not as a patch to integrate upstream.
Further optimizations of the internal CUDACPP engine for computing matrix elements, possibly extending the work I presented here, are of course possible. The clean separation of Feynman diagrams and color sums in this new implementation, for instance, makes it possible to imagine scenarios where the former is performed on vector CPUs and the latter on GPUs using cuBLAS for particularly complex processes. In my opinion, however, the ME calculation for LO processes in CUDACPP has now been optimized to a level where there are no obvious opportunities for further significant throughput speedups. Other areas of work have higher priority, such as in particular the integration of LO CUDACPP into the upcoming Madgraph7 with an enhanced phase space sampling based on MadNIS bib:madnis , or the extension of the CUDACPP infrastructure to NLO physics processes bib:zwnlo . The work I presented in this paper will be relevant in both of these activities.
Acknowledgements
I thank the MG5aMC and the Madgraph on GPU development teams for our fruitful collaboration on these projects. In particular, I warmly thank Olivier Mattelaer for our initial discussions and our coding together for the new developments shown in this paper during the 2022 GPU Hackathon and more generally for our effective collaboration over many years on the Madgraph on GPU developments. This project would not have been possible without him. I thank Daniele Massaro for useful discussions about SIMD support in CUDACPP and for his help in the integration of this work in new production releases of the CUDACPP software. I warmly thank Domenico Giordano for many useful discussions about hardware benchmarking. I am grateful to the organisers and our mentors at the CSCS GPU Hackathon in Lugano in September 2022, and I thank the other participants from our team, Stephan Hageböck, Stefan Roiser and Zenny Wettersten, for our useful discussions. I kindly acknowledge the use of LUMI HPC resources under project 465001592 (CERN / Madgraph GPU porting, EHPC-DEV-2024D11-007) to derive the results I showed for AMD GPUs.